### v1

# Equation-of-state-informed pulse profile modeling

Abstract
NICER has enabled mass–radius inferences for pulsars using pulse profile modeling, providing constraints on the equation of state (EOS) of cold, dense matter. To date, PPM and EOS inference have been carried out as two separate steps, with the former using EOS-agnostic priors. This approach has several drawbacks. Ideally, one would perform a fully hierarchical Bayesian inference where the pulse profile and EOS model parameters are jointly fit, but implementing such a framework is complex and computationally demanding. Here, we present an intermediate solution introducing an EOS-informed prior on mass-radius into the existing PPM pipeline using normalizing flows. By focusing on the parameter space consistent with certain EOSs, this approach both tightens constraints on neutron star parameters while reducing computational costs and requiring minimal additional implementation effort. We test this approach on two pulsars, PSR J0740+6620 and PSR J0437-4715, and with two EOS model families: a model based on the speed of sound inside the neutron star interior (CS) and a piecewise-polytropic (PP) model. Both EOS models implement constraints from chiral effective field theory calculations of dense matter. For both pulsar datasets, the inferred radius credible intervals are narrower than in the EOS-agnostic case, with CS favoring smaller radii and PP favoring larger radii. For PSR J0437-4715, the EOS-informed priors reveal a new, more extreme geometric mode that is statistically favored but physically questionable. Including the PPM posteriors in the subsequent EOS inference further tightens the mass-radius posteriors through the chiral effective field theory constraints. However, there is also a sensitivity to the high-density extensions, where the PP (CS) model produces a shift towards larger (smaller) radii and corresponding stiffening (softening) of the pressure–energy density relation.
I Introduction
Neutron stars, with their extreme compactness, provide a unique opportunity to study the behavior of dense nuclear matter and constrain the equation of state (EOS) (see, e.g., Watts et al., 2016; Lattimer and Prakash, 2016; Baym et al., 2018; Tolos and Fabbietti, 2020). Observations of X-ray emission from their surfaces serve as a powerful tool for probing their internal structure, thereby placing constraints on the EOS of cold, dense matter.
The Neutron Star Interior Composition Explorer (NICER; Gendreau et al. (2016)), onboard the International Space Station, has played a pivotal role in this effort by detecting soft thermal X-rays from a specific class of neutron stars: rotation-powered millisecond pulsars. These pulsars exhibit rotationally modulated, pulsed X-ray emission, believed to originate from hot spots at the magnetic poles where return currents deposit heat (see, e.g., Ruderman and Sutherland, 1975; Arons, 1981; Harding and Muslimov, 2001). The X-ray signal is analyzed using the pulse profile modeling (PPM) technique (e.g., Watts, 2019; Bogdanov et al., 2019, 2021, and references therein), which incorporates relativistic effects arising from the neutron star’s rapid spin and intense gravitational field. This approach enables precise measurements of neutron star masses and radii.
To date, NICER has enabled mass and radius inferences for five pulsars: PSR J0030+0451, Riley et al. (2019); Miller et al. (2019); Salmi et al. (2023); Vinciguerra et al. (2024), PSR J0437-4715 Choudhury et al. (2024), PSR J0740+6620 Miller et al. (2021); Riley et al. (2021); Salmi et al. (2022, 2023, 2024a); Dittmann et al. (2024); Hoogkamer et al. (2025), PSR J1231-1411 Salmi et al. (2024b); Qi et al. (2025), and PSR J0614-3329 Mauviard et al. (2025). The resulting mass-radius (MMRR) constraints have been widely used to constrain the EOS of cold, dense matter (see, e.g., Rutherford et al., 2024; Koehn et al., 2024; Huang et al., 2025; Golomb et al., 2025; Li et al., 2025).
The X-ray Pulse Simulation and Inference (X-PSI; Riley et al., 2023) code is an open-source software package for PPM and Bayesian statistical inference. [^1] https://github.com/xpsi-group/xpsi In MRM–R inference to date, X-PSI has adopted an EOS-agnostic approach, following the methodology in, e.g., Riley et al. (2018); Raaijmakers et al. (2018), selecting joint flat priors on mass and radius, and modifying the mass prior when informative radio-timing measurements are available. EOS inference, using the open-source software NEoST Raaijmakers et al. (2025), [^2] https://github.com/xpsi-group/neost is then performed as a separate analysis step. The main motivation for this two-step approach is to provide the community with posterior samples in the MMRR plane that can be studied using any EOS model afterwards. However, this strategy has some practical drawbacks:
  1. 1.
    High computational costs: the accuracy and precision of the PPM-inferred neutron star properties depend on factors such as data quality and the availability of prior information. When these are limited, the resulting parameter space becomes broad or difficult to explore efficiently. For instance, inference for PSR J0437-4715 required $\sim 783$k core hours to achieve convergence Choudhury et al. (2024).
  2. 2.
    Sampling of implausible parameter space: the use of flat priors permits sampling of regions, such as very compact solutions, that are already ruled out by physical EOS constraints - yet these regions incur very expensive multiple-image ray-tracing computations.
  3. 3.
    Difficulty with multi-modality: a source like PSR J0030+0451 Salmi et al. (2023); Vinciguerra et al. (2024) exhibits complex, multimodal solutions that are difficult to sample thoroughly within reasonable compute time with flat priors. Imposing realistic priors that exclude regions already excluded by EOS constraints will focus sampling only on physically meaningful regions of parameter space, and potentially break multi-modality and reduce sampling time.
  4. 4.
    Because $M$ and $R$ are generally correlated with geometry parameters (e.g., spot size and location), imposing more physically meaningful priors will also lead to inferred geometries that are more realistic and better constrained, in contrast to the current implementation.
In an ideal scenario, one would perform fully hierarchical Bayesian inference where the pulse profile and EOS parameters are jointly fit. While this would solve the aforementioned issues, the implementation of such a framework is very complex, time-consuming, and running it could be computationally intractable.
At present, the X-PSI and NEoST packages are two distinct frameworks, each focusing on different aspects of the full Bayesian model, namely PPM and EOS inference, respectively. PPM is carried out for individual stars inferring parameters such as mass and radius. EOS parameters are introduced only in the second stage, along with stellar central densities, when information from multiple stars is combined. Properly unifying the PPM and EOS inference would substantially increase the number of parameters, as it would require simultaneous inference of neutron star properties, EOS parameters, and central density even for a single star. This framework would eventually make it possible to carry out PPM for multiple stars at the same time. On the other hand, such a joint inference could benefit from Bayesian shrinkage: the requirement that all stars share the same EOS may help break certain parameter degeneracies, potentially making the overall sampling process more efficient.
As an intermediate step, in this work, we introduce an EOS-informed MMRR prior into the existing PPM pipeline, while leaving the subsequent EOS inference unchanged. This approach requires minimal additional implementation effort but still constrains the PPM step to regions of parameter space consistent with prior EOS information, thereby providing a way to obtain tighter constraints on neutron star parameters through more physically informed priors. In addition, it helps reduce computational costs by limiting the parameter space that must be sampled, which in turn might lead to quicker convergence. Finally, leaving the PPM and EOS inference as two separate steps has the additional advantage of avoiding having to deal with a substantial increase in the number of parameters as would be the case when unifying the two.
To inform the PPM pipeline of the underlying cold, dense matter EOS prior, we use the two EOS model families considered in Rutherford et al. (2024). At nuclear densities beyond the BPS crust (Baym et al., 1971), these are based on next-to-next-to-next-to-leading-order (N3LO) chiral effective theory calculations of matter in beta equilibrium from Keller et al. (2023) up to $1.5n_{0}$, where $n_{0}=0.16\,\mathrm{fm}^{-3}$ is the nuclear saturation density. The EOS-prior thus builds in knowledge of the crust and of the EOS at nuclear densities based on our modern understanding of nuclear forces. At higher densities, we use two possible model extensions, based on the speed of sound inside the neutron star interior (CS) Greif et al. (2019) or based on a piecewise-polytropic (PP) model Hebeler et al. (2013). In this work, we explore the impact of both prior EOS ensembles on the PPM-derived MMRR and geometry posteriors of PSR J0740+6620 and PSR J0437-4715, and the effects of the derived MMRR posteriors on the subsequent EOS inference analysis.
The remainder of this paper is organized as follows. In Sec. II.1, we present the ideal case of a full hierarchical Bayesian model for the simultaneous inference of neutron star and EOS properties. Section II.2 introduces our intermediate approach that incorporates EOS prior information to constrain neutron star properties, followed by the EOS inference framework using EOS-prior-informed PPM MRM–R posteriors. Sections II.3 and II.4 describe the X-PSI PPM analysis pipeline and the use of normalizing flows to model the MRM–R prior derived from the two EOS families. In Sec. III, we summarize the datasets for PSR J0740+6620 and PSR J0437–4715, and outline the prior intervals for the PP and CS EOS model parameters used in this analysis. Section IV presents the EOS-prior-informed PPM posteriors and the resulting constraints on the cold, dense matter EOS. Finally, in Sec. V, we discuss our conclusions and directions for future work.
II Methodology
Section II.1 outlines the ideal scenario of a full hierarchical Bayesian model for the joint inference of neutron star and EOS parameters. In Sec. II.2, we introduce our intermediate method—that does not require substantial implementation effort—which incorporates EOS prior information to constrain neutron star properties and enables EOS inference based on EOS-prior-informed PPM MRM–R posteriors. Sections II.3 and II.4 detail the X-PSI PPM analysis pipeline and the implementation of normalizing flows to represent the MRM–R prior derived from the two EOS models, respectively. Complete information of each PPM and EOS inference run, including the exact X-PSI and NEoST versions, data products, posterior sample files, and all the analysis files are available in a Zenodo repository. [^3] https://doi.org/10.5281/zenodo.17257403
II.1 Full Hierarchical Bayesian Model
Currently, posterior densities are derived for the parameters of each neutron star independently, and later, the $M$-$R$ posteriors are used as “data” in EOS inference. However, under the assumption that all neutron stars adhere to the same EOS, it is much more informative to infer both global EOS parameters and the local parameters for all available neutron stars simultaneously, using a hierarchical Bayesian model. A visual representation of this full model, using probabilistic graphical models, is shown in Fig. 1. A key advantage of this approach is the occurrence of Bayesian shrinkage: by constraining all neutron stars to lie on a common EOS curve, we should obtain tighter constraints on the MRM–R relation and the EOS parameters.
We begin by representing the pulse profile data as $\left\{D_{k,j}\right\}^{K,J}_{k,j=1}$, where $D_{k,j}$ denotes the observed data for the $k$-th neutron star with $K$ the total number of neutron stars included in the analysis, and the $j$-th data point with $J$ the total number of data points in each pulse profile. See Tab. 1 for an overview of all parameter definitions.
Figure 1: Probabilistic graphical model illustrating the hierarchical Bayesian framework for the joint inference of neutron star properties and EOS parameters. Arrows denote statistical dependencies between variables. The shaded gray circle represent observed variables, while double circles indicate deterministic variables. The box represents a summation that applies to all parameters contained within it. See Tab. 1 for a detailed description of all variables.
Next, we specify the prior distribution for the parameters not directly affected by the EOS, such as the inclination angle and the hot spot geometry parameters:
$p(\left\{\theta_{k}\right\}^{K}_{k=1})=\prod_{k=1}^{K}p(\theta_{k})\,$, (1)
with the parameter $\theta_{k}$ not affected by the EOS, assuming independence between different neutron star sources.
Parameter Description
$\left\{D_{k,j}\right\}^{K,J}_{k,j=1}$
Observed pulse profile data for neutron stars
k=1,,Kk=1,...,K for each data point j=1,,Jj=1,...,J
$\left\{\theta_{k}\right\}^{K}_{k=1}$ Parameters not affected by the EOS
$\left\{M_{k}\right\}^{K}_{k=1}$ Gravitational mass for each neutron star $k$
$\left\{R_{k}\right\}^{K}_{k=1}$ Equatorial radius for each neutron star $k$
$\varphi$ EOS parameters
$\left\{\varepsilon_{c,k}\right\}^{K}_{k=1}$ Central energy density for each neutron star $k$
Table 1: Definitions for each parameter of the statistical model.
In contrast, the neutron star parameters that are affected by the EOS are governed by a deterministic relationship. Using ($\varphi,\varepsilon_{c,k}$), the vector containing the EOS parameters and the central density of the $k$-th neutron star, respectively, and solving the Tolman-Oppenheimer-Volkoff (TOV; see Tolman, 1939; Oppenheimer and Volkoff, 1939) equations, the radius $R_{k}$ is determined by the MMRR relationship of the EOS for a given sample from the prior for $M_{k}$. $M_{k}$ is sampled from a mass prior based on radio timing observations (when available), or from a uniform prior (when such observations are unavailable), and $R_{k}$ is determined using $\varphi$, $\varepsilon_{c,k}$ and $M_{k}$. In this setup, the mass $M_{k}$ is treated as an independent variable, and we can express its prior as
$p(\left\{M_{k}\right\}^{K}_{k=1})=\prod_{k=1}^{K}p(M_{k})\,$, (2)
assuming independence between different neutron stars.
The prior on the EOS parameters $\varphi$ is given by
$p(\varphi)\,.$ (3)
Similarly, the prior on the central energy density $\varepsilon_{c,k}$ is given by:
$p(\left\{\varepsilon_{c,k}\right\}^{K}_{k=1})=\prod_{k=1}^{K}p(\varepsilon_{c,k})\,$, (4)
again assuming independence across neutron stars. We note that, strictly speaking, $\left\{\theta_{k}\right\}^{K}_{k=1}$, $\left\{\varepsilon_{c,k}\right\}^{K}_{k=1}$, and $\varphi$ depend on certain fixed hyperparameters. [^4] Hyperparameters are parameters that determine the shape of the prior distribution for a given model parameter. For example, if the prior for an inclination angle follows a normal distribution, the hyperparameters would be its mean and standard deviation. However, since these hyperparameters are fixed and therefore not random variables, we have chosen not to write them down for clarity. Similarly, all distributions in the joint PPM-EOS inference should be read as conditional on the choices of PPM model (see, e.g., Riley et al., 2019; Vinciguerra et al., 2023) and on the choice of EOS model (see Hebeler et al., 2013; Greif et al., 2019, for the PP and CS models, respectively). To avoid cluttering the notation, we have omitted including these in the conditional densities as well.
The likelihood is described by
[…]
where we assume that both neutron stars and data points within each pulse profile are statistically independent. By combining Eqs. (1), (2), (3), (4), and (5), we obtain the posterior probability distribution:
$p(\left\{\theta_{k},M_{k},\varepsilon_{c,k}\right\}^{K}_{k=1},\varphi|\left\{D_{k,j}\right\}^{K,J}_{k,j=1})$
$\propto\prod_{k=1}^{K}\left[\prod_{j=1}^{J}p(D_{k,j}|\theta_{k},M_{k},\varepsilon_{c,k},\varphi)\right]$
$\times~p(\theta_{k})p(M_{k})p(\varepsilon_{c,k})p(\varphi)\,.$ (6)
As mentioned before, since the equatorial radius $R_{k}$ of a given source is completely determined by the combination of $(\varphi,\varepsilon_{c,k},M_{k})$, the posterior distribution on $R_{k}$ can be computed once the posteriors on $\varphi$, $\varepsilon_{c,k}$, and $M_{k}$, are determined.
II.2 Intermediate Step: PPM including EOS Prior Information and EOS Inferencing
[…]
To construct the EOS-informed posterior distribution for the derived neutron star properties, we begin by formulating the posterior obtained via X-PSI. In previous X-PSI analyses of rotation-powered millisecond pulsars (e.g., Riley et al., 2019; Salmi et al., 2023; Vinciguerra et al., 2024; Riley et al., 2021; Salmi et al., 2022, 2023, 2024a; Hoogkamer et al., 2025; Choudhury et al., 2024), analysis was done for a single star ($K=1$) and the EOS prior was not incorporated. Instead, the posterior was given by:
$p(\theta,R,M|\left\{D_{j}\right\}^{J}_{j=1})\propto\prod_{j=1}^{J}p(D_{j}|\theta,R,M)p({\theta)p({R)p(M)}}\,.$ (7)
where no explicit EOS model was used to constrain the MRM–R relation.
To fold in the EOS prior information into the MMRR space within X-PSI, we modify Eq. (7) to define a new posterior:
$p\big($ $\theta,R,M|\left\{D_{j}\right\}^{J}_{j=1},\varphi\big)$
$\propto\prod_{j=1}^{J}p(D_{j}|\theta,\varphi,R,M)p(\theta)p\big(R|\varphi\big)p\big(M\big)p\big(\varphi\big)\,$, (8)
where the EOS parameters $\varphi$ now inform the MMRR distribution. In practice, we adopt the following procedure. First, we sample the priors on the EOS parameters and compute their corresponding priors in the MMRR space using the NEoST framework (described in more detail below), as shown in Fig. 2. We then use these samples to train a normalizing flow (see Sec. II.4) that approximates the EOS MMRR prior. Subsequently, if no mass prior is available from radio timing, we draw the mass $M$ from a flat prior. If a mass prior from radio timing is available, we instead sample a mass $M$ from that prior. In either case, we then draw a corresponding radius sample $R$ from the EOS $M$-$R$ prior conditional on the sampled mass $M$.
[…]
For the EOS inference, we implement the Raaijmakers et al. (2021) Bayesian framework using NEoST (v2.1.0)Raaijmakers et al. (2025). [^5] The pre-release form of NEoST was used in (Greif et al., 2019; Raaijmakers et al., 2019, 2020, 2021; Rutherford et al., 2023). The released version of NEoST has been used in Rutherford et al. (2024, 2025). NEoST uses MultiNest to sample from the chosen EOS parameter space and combines constraints from both the MRM–R relations derived from PPM and the $M–\Lambda$ relations from gravitational-wave observations. In this work, however, we focus solely on incorporating EOS prior information into the PPM-derived MRM–R posteriors, and therefore omit the gravitational-wave constraints to highlight the differences between the original and new frameworks. Since the EOS model parameters considered in this work—namely PP and CS—are uniformly sampled, the EOS inference framework is identical to that used in our previous work Raaijmakers et al. (2021); Rutherford et al. (2024).
The two-step process implemented here naturally raises concerns about double-counting the EOS prior, since it is used both to define priors for MMRR during the PPM stage and again in the subsequent EOS inference. This is especially relevant because uniform priors on the EOS parameters lead to non-uniform priors on neutron star mass and radius due to the non-linearity of the TOV equations. However, since these equations are deterministic, the transition from EOS parameters to MMRR space corresponds to a change of variables, which is a common operation in Bayesian inference especially when it reduces the computational costs of sampling. Since in our case the prior on the EOS parameters is flat, applying them during both the PPM and subsequently the EOS inference does not distort the overall shape of the EOS posterior, and thus does not influence EOS parameter estimation. It is important to note, however, that Bayesian model comparison via posterior odds should not be performed in this manner, as the inclusion of the prior in both inference steps would indeed distort the results.
Therefore, the posterior distribution of the EOS parameters $\varphi$ and central densities $\varepsilon_{c}$ remains unchanged and is given by
$p(\bm{\varphi},\bm{\varepsilon_{c,k}}\,|\,\bm{d})\propto p(\bm{\varphi})p(\bm{\varepsilon_{c,k}}\,|\,\bm{\varphi})\prod_{k=1}^{K}p(M_{k},R_{k}\,|\,\bm{d}_{k})\,$, (9)
where $\bm{d}$ is the dataset used to constrain the EOS. In this work, $\bm{d}$ consists of the MRM–R PPM posteriors for PSR J0740+6620 and PSR J0437-4715, obtained using an EOS-informed MRM–R prior together with a mass prior derived from radio timing. We do not include other data, such as gravitational wave observations, in the EOS inference.
It is important to note that when an EOS-informed prior is adopted in the PPM step, the EOS inference is only self-consistent if the same EOS model is used in both PPM and EOS inference. EOS inference with a different EOS model family than that used to define MMRR priors for PPM is still possible, provided that the EOS model used to define the PPM priors encompasses the full parameter range of the EOS model applied during the EOS inference. In this scenario, EOS inference is akin to using PPM results derived using EOS-agnostic priors. [^6] It is important to note that applying different EOS models between the PPM and EOS analysis stages is dependent on whether or not the priors on the EOS model used in the later stage are uniform. If the priors for the EOS model parameters are flat, similar to what is done in this work, then the EOS inference will not be affected by the shape of those priors. However, if the EOS priors are non-uniform, then there is an effective double counting of the prior contribution of the EOS model, thus making the EOS posterior inference statistically inconsistent. This ensures that the prior informing the PPM-derived MRM–R distributions—the EOS parameters $\varphi$—remains aligned with the model used in the subsequent EOS inference.
II.3 Pulse Profile Modeling using X-PSI
We use the open-source X-PSI package, with versions ranging from v3.0.0 to v3.0.5 Riley et al. (2023). As done previously (see e.g., Riley et al., 2021; Salmi et al., 2022, 2024a; Choudhury et al., 2024), this uses the ‘Oblate Schwarzschild + Doppler’ approximation to model the energy-resolved X-ray pulses from the neutron star Miller and Lamb (1998); Nath et al. (2002); Poutanen and Gierliński (2003); Morsink et al. (2007); Lo et al. (2013); AlGendy and Morsink (2014); Bogdanov et al. (2019); Watts (2019). This approximation takes the oblate shape of the neutron star into account and the relativistic effects resulting from the rapid spin of the rotation-powered millisecond pulsars while treating the exterior spacetime as Schwarzschild. For a more detailed description of X-PSI, see Riley et al. (2019, 2023).
X-PSI is used in combination with the nested sampling algorithm MultiNest Feroz and Hobson (2008); Feroz et al. (2009, 2019) and its Python binding PyMultiNest Buchner (2016) to compute posterior samples and calculate the evidence for Bayesian model comparison. We opted for this set-up over UltraNest Buchner (2021) due to its lower computational cost, despite UltraNest’s proven convergence and stable results for PSR J0740+6620 Hoogkamer et al. (2025).
For the massive pulsar PSR J0740+6620, we use the exact same setup and sampler settings as in the headline result of Salmi et al. (2024a) and Hoogkamer et al. (2025), where NICER and XMM-Newton data are jointly modeled. This setup includes informative priors on mass, inclination, and distance obtained from radio timing Fonseca et al. (2021). Section II.4 describes how this mass prior is combined with the EOS-informed MMRR prior.
For the nearby and bright pulsar PSR J0437–4715, we likewise adopt the same setup as in the headline result of Choudhury et al. (2024). This setup incorporates tight informative priors on mass, inclination, and distance based on radio timing Reardon et al. (2024) (see Sec. II.4 for more details), as well as NICER background models to constrain the non-source background, cross-validated with XMM-Newton data. Instead of 20k live points and a sampling efficiency of 0.3 used by Choudhury et al. (2024), we use 4k live points and 0.1 sampling efficiency for the EOS-informed prior runs to save on computational costs. These settings are expected to be sufficient, as testing the EOS-agnostic case with 4k live points and 0.1 sampling efficiency yielded results consistent with those obtained using 20k live points and 0.3 sampling efficiency in Choudhury et al. (2024).
II.4 Normalizing Flow as EOS Prior
Uniform sampling in the EOS space induces a non-uniform prior in the MMRR space due to the non-linearity of the TOV equations. Because an analytical transformation of the probability density between these spaces is intractable, we approximate the MMRR prior to enable sampling from it during the PPM stage. To achieve this, we draw samples from the uniform prior in EOS space, transform them into corresponding pairs in MMRR space by solving the TOV equations, and then fit a flexible parametric form to the resulting distribution using a normalizing flow.
Normalizing flows (see, e.g., Dinh et al., 2014; Papamakarios et al., 2019; Kobyzev et al., 2019) are a class of deep learning models used for density estimation, i.e., learning a function that approximates an unknown probability distribution. Normalizing flows approximate the target distribution by transforming a simple and analytically known distribution, such as a standard normal distribution, into a more complex one by applying a sequence of invertible (i.e., bijective) and differentiable transformations. These transformations are parameterized by neural networks to allow for greater flexibility, and enable both sampling and evaluation of the log-density of the resulting distribution.
In this work, we use normalizing flows to model the EOS MMRR prior distribution,
p(M,R)=p(R|M)p(M),p(M,R)=p(R|M)p(M)\,, (10)
allowing us to efficiently sample joint MMRR combinations, as well as draw radius samples conditioned on a given mass. The latter is particularly useful in cases where a mass prior is available from radio timing observations, as for PSR J0740+6620 and PSR J0437-4715.
In this paper, we use the implementation of normalizing flows provided in Pyro Bingham et al. (2018), [^7] https://pyro.ai with a standard normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=1)$ as the base distribution. For both mass and radius, we use transformations known as coupling neural spline flows Durkan et al. (2019); Dolatabadi et al. (2020); Dinh et al. (2014); Müller et al. (2018), specifically element-wise rational-linear spline bijections, where the radius is conditioned on the mass. Each transformation is parametrized by a small dense feedforward neural network that takes MMRR samples from the EOS prior distributions (see Sec. III.3) as input. The dataset consists of 100,000 MMRR samples of which we reserve 70% for training, 15% for validation, and another 15% for testing. Prior to training, the data are standardized using Pyro’s StandardScaler to have zero mean and unit variance.
To determine the architecture and training hyperparameters, [^8] In this context, hyperparameters are settings chosen a priori that control how the neural network learns, such as the learning rate and the number of hidden layers. we perform Bayesian hyperparameter tuning using W&B Biewald (2020), [^9] https://wandb.ai aiming to select a configuration that balances validation loss (the negative log-likelihood of the samples given the transformation of the distribution) with training efficiency. The sweep varied the hidden layer dimensions (8408~–~40), the number of transformations (1101~–~10) of $p(R|M)$ and $p(M)$ as shown in Eq. (10), the learning rate ($5\times 10^{-4}~-~5\times 10^{-2}$), and the number of training epochs (25012000250~–~12000 in steps of 250).
After performing the Bayesian hyperparameter tuning, we selected the best-performing configuration without further exploring alternative setups. Because there were many configurations with similar validation losses, we chose one with a relatively short training time. This configuration—using a normalizing flow with ten transformations for both $p(R|M)$ and p(M)p(M)—performed well, achieving a test loss of 2.52 for CS and 2.14 for PP (see Fig. 3 for the training and validation loss curves) in approximately 27 minutes for both cases. Other configurations with comparable loss values required substantially longer training times. The resulting network contains two hidden layers with 32 neurons each, and ReLU [^10] ReLU (Rectified Linear Unit) is an activation function defined as $f(x)=\text{max}(0,x)$. Activation functions introduce non-linearity into the network, allowing the model to capture complex patterns and relationships in the data. Nair and Hinton (2010) activations following each layer. The normalizing flow is trained on 70,000 MRM–R samples for 4000 epochs, using the Adam optimizer Kingma and Ba (2017) with a learning rate of $3\times 10^{-3}$. The model is optimized by minimizing the negative log-likelihood of the MMRR samples.
Figure 3: Training (orange) and validation (blue) loss curves for the PP EOS model (top) and CS EOS model (bottom).
To evaluate, we compare the probability distribution learned by the normalizing flow to our test set using the Kullback-Leibler (KL) divergence, which provides a single number quantifying how closely the entire learned distribution approximates the target one. This results in KL-divergences of 0.0075 and 0.0023 for the normalizing flow of the CS and PP EOS prior distribution, respectively. A KL-divergence close to zero indicates that the distribution learned by the normalizing flow is nearly indistinguishable from the test set. See Fig. 2 for a visual comparison between the prior model and the resultant normalizing flow.
When sampling radii conditioned on a given mass, as in the cases of PSR J0740+6620 and PSR J0437-4715, samples with radii outside the interval 5.5 to 14.0 km for the PP EOS model and 6.5 to 14.5 km for the CS EOS model are excluded from the analysis. These bounds are chosen to avoid artifacts arising in regions where the normalizing flow assigns low but nonzero probability and to avoid drawing unphysical MMRR values. The bounds are conservative and lie well outside the support of the EOS prior. For a sample size of one million, the probability of drawing samples outside these bounds is close to zero, so truncation is not expected to significantly affect the normalization of the distribution.
III Data
In this section, we outline the X-ray event data used in this study. In Sec. III.1, we describe the most important aspects of the dataset from the high-mass pulsar PSR J0740+6620, as detailed in Salmi et al. (2024a); Hoogkamer et al. (2025). Section III.2 gives a brief description of the observations from the nearest and brightest rotation-powered millisecond pulsar PSR J0437-4715, as detailed in Choudhury et al. (2024). In Sec. III.3, we summarize the prior distributions of the PP and CS EOS models.
III.1 PSR J0740+6620
For PSR J0740+6620, we use the NICER dataset between September 21, 2018, and April 21, 2022, identical to Salmi et al. (2024a); Hoogkamer et al. (2025). We use the pulse-invariant channel subset [30, 150), corresponding to the nominal photon energy range [0.3, 1.5] keV. The data are processed in the same way and include 2.73 Ms (roughly a month) of on-source exposure time after filtering. For background constraints, the same XMM-Newton phase-averaged spectral data and blank-sky observations are used with the three EPIC instruments (pn, MOS1, MOS2), including the energy channels [57, 299) for pn, and [20, 100) for both MOS1 and MOS2 (as reported in Riley et al. (2021); Wolff et al. (2021); Salmi et al. (2022, 2024a); Hoogkamer et al. (2025)). Furthermore, the hot emitting regions of the neutron star are modeled with two independent circular uniform temperature spots using the ST-U (Single Temperature-Unshared) model (see Riley et al. (2019); Vinciguerra et al. (2023) for more details on the X-PSI model naming convention and parameters).
III.2 PSR J0437–4715
Similar to the headline result of PSR J0437-4715 in Choudhury et al. (2024), we use the 3C50 NICER dataset, which contains observations from July 6, 2017 to October 11, 2021. This dataset owes its name to the usage of the 3C50 background model, which is designed to estimate the instrumental background in NICER observations. The model utilizes NICER event data to predict the background level as a function of energy (see Remillard et al. (2022) for the model definition). To get the best estimate of the instrumental background of NICER, the data are filtered according to specific criteria (see Salmi et al. (2022) for a detailed description of the analysis and filtering procedure). The filtering steps reduce the usable exposure time from 2.736 Ms to 1.328 Ms, but enable the quantification of the uncertainties in the net background spectrum, which is predominantly influenced by systematic instrumental effects.
To maximize the signal-to-noise ratio and mitigate contamination from the bright Seyfert II active galactic nucleus RX J0437-4711, located $4^{\prime}.18$ from PSR J0437-4715, NICER observations were conducted pointing $1^{\prime}.15$ south-west of the pulsar. This deliberate offset reduces the AGN flux within the field of view to only $\sim$2.25% (i.e., $\sim 0.2$ counts/s). The AGN flux is then modeled as an additive background component.
For the analysis, only the events in the pulse-invariant channels [30, 300) corresponding to the nominal photon energy range 0.3–30 keV are kept. This is done to match the energy range from XMM-Newton observations, which are used to cross-validate the results from NICER. Lastly, the hot emitting regions of the neutron star are modeled with a two-temperature spot and a ring using the CST+PDT (Concentric Single Temperature + Protruding Dual Temperature model). The two-temperature hot spot consists of two overlapping spherical emitting regions, referred to as the ceding and superseding components. The ring is modeled with two spherical components with one emitting and one masking (the latter referred to as omitting in Vinciguerra et al. (2023, 2024)).
III.3 PP and CS EOS Prior Distributions
With the EOS-informed PPM framework in place (Sec. II.2), the next step is to compute the MMRR priors for the PP and CS EOS models. This involves constructing the EOS models across the relevant density range and defining the prior ranges that specify their distributions. Following Rutherford et al. (2024), we briefly summarize the EOS construction adopted in this work and the corresponding priors.
[…]
For the PP model, the priors are defined by varying three polytropic indices, $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$, as well as two transition densities, $n_{12}$ and $n_{23}$, which connect the first to the second polytrope and the second to the third, respectively. The priors on these parameters are given by: $\Gamma_{1}\in[0,8]$, $\Gamma_{2}\in[0,8]$ $\Gamma_{3}\in[0.5,8]$, and $2n_{0}\leq n_{11}\leq n_{23}\leq 8.3n_{0}$.
For the CS model, the prior distribution is defined by parameterizing the speed of sound $c_{s}^{2}=dP/d\varepsilon$, using five parameters $a_{1}$ to $a_{5}$, which are varied over the following ranges: $a_{1}\in[0.1,1.5]$, $a_{2}\in[1.5,12]$, $a_{3}/a_{2}\in[0.05,2]$, $a_{4}\in[1.5,37]$, and $a_{5}\in[0.1,1]$. The resulting prior distributions for both the PP and CS models span a large range of possible EOSs. For additional details on the PP and CS EOS models, see Rutherford et al. (2024); Hebeler et al. (2013); Greif et al. (2019)
Finally, all EOS priors and posteriors are required to obey causality and the constraint that neutron stars have masses $\geq 1 MM_{\odot}$. This lower mass limit is imposed to be consistent with the lowest confirmed neutron star mass observation Martinez et al. (2015) and is also motivated by theoretical models of early neutron star evolution Strobel et al. (1999) as well as the minimum remnant masses predicted by simulations of core-collapse supernovae Janka et al. (2008); Fischer et al. (2010); Radice et al. (2017); Suwa et al. (2018); Müller et al. (2025). For completeness, the resulting MMRR prior distributions of both of the PP and CS models are shown in Fig. 2.
IV Results
The results, aimed at showing the impact of incorporating the EOS prior information in the analysis, are presented in two parts. In Sec. IV.1, we show the PPM inference results using the EOS MMRR prior information for both PSR J0740+6620 and PSR J0437-4715. In Sec. IV.2, we present the results of the subsequent EOS inference using the posteriors from Sec. IV.1.
IV.1 PPM including EOS Prior Information
To assess the influence of an EOS-informed MRM–R prior on the inferred properties of PSR J0740+6620 and PSR J0437-4715, we evaluate two representative EOS models as priors: CS and PP. These are compared against the baseline EOS-agnostic approach.
IV.1.1 PSR J0740+6620
For PSR J0740+6620, we take the headline results of Salmi et al. (2024a), using the EOS-agnostic approach, as our baseline. The analysis of Salmi et al. (2024a), using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01, reported an equatorial radius of $12.49^{+1.28}_{-0.88}$ km and a gravitational mass of $2.073^{+0.069}_{-0.069}$ $M_{\odot}$. Figure 4 shows the posterior distributions, as well as the ones obtained when incorporating the CS and PP models as prior information on MRM–R. Tab. 2 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior for each model.
Figure 4: Radius, compactness, and mass posterior distributions using the PSR J0740+6620 joint NICER and XMM-Newton dataset conditional on the ST-U model. Three posterior distributions are shown: the results from Salmi et al. (2024a) without an EOS-informed prior (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the PP-informed MMRR prior (blue). All inference results are obtained using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01. The marginal prior probability distribution functions for each parameter are displayed as dashed-dotted lines. The shaded regions in the diagonal panels contain the 68.3% credible interval for each parameter symmetric around the median. The contours in the off-diagonal panels contain the 68.3%, 95.4%, and 99.7% credible regions.
The radius credible intervals for all three models overlap. However, the radius inferred using the CS-informed prior is $11.90^{+0.60}_{-0.55}$ km—with the median smaller than the baseline—while the PP-informed prior yielded a slightly larger median radius of $12.52^{+0.54}_{-0.61}$ km. The results are in line with our expectation, since the CS model favors softer EOSs and the PP model favors stiffer ones. However, once several sources are included, the CS and PP differences are smaller Rutherford et al. (2024). Notably, the mass and radius credible intervals obtained with the EOS-informed priors are narrower than those obtained using the EOS-agnostic MMRR prior.
The posteriors of the geometry parameters, as shown in Fig. 10 of the Appendix, match even better for both the EOS-agnostic approach and the EOS-informed prior runs, with overlapping credible intervals for all parameters. Furthermore, the difference between the maximum likelihood and Bayesian evidence between all three models is negligible. [^11] The difference in maximum log-likelihood is approximately $0.0$ across all three cases, while the evidence differs by approximately +0.30+0.30 for the CS-informed run and +0.36+0.36 for the PP-informed run relative to the EOS-agnostic approach.
In terms of computational cost, the EOS-agnostic inference required 84k core hours, compared to 41k and 36k core hours for the runs using the CS-informed and PP-informed priors, respectively.
Parameter
Salmi et al. 2024
(EOS agnostic)
CS-informed
$M$-$R$ prior
$P$-informed
$M$-$R$ prior
$R_{\text{eq}}[\text{km}]$ 0.64 0.38 0.15
$M/R_{\text{eq}}$ 0.72 0.54 0.16
$MM_{\odot}$]] 0.01 0.03 0.01
Table 2: KL divergence (in bits) between the posterior and corresponding prior for each parameter, computed for the three different inferences for PSR J0740+6620 shown in Fig. 4.
IV.1.2 PSR J0437-4715
For PSR J0437–4715, we adopted the headline result from Choudhury et al. (2024) as our baseline, using MultiNest with 20k live points and a sampling efficiency of 0.3. For the EOS-informed prior runs, we used $4\times 10^{3}$ live points and a sampling efficiency of 0.1 to reduce computational costs. The resulting posterior distributions for all three models are shown in Fig. 5. Tab. 3 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior.
As shown in Fig. 5, the EOS-agnostic approach yielded an equatorial radius of $11.36^{+0.95}_{-0.63}$ km and a gravitational mass of $1.418^{+0.037}_{-0.037}$ $M_{\odot}$. The CS-informed prior run inferred a radius of $11.03^{+0.43}_{-0.41}$ km, which is slightly smaller than the baseline result, though the credible intervals otherwise strongly overlap. On the other hand, the PP-informed prior gives a radius of $12.92^{+0.25}_{-0.28}$ km, which is noticeably larger. Similar to PSR J0740+6620, these results align with expectations: the CS model favors softer EOSs given a smaller radius, while the PP model favors stiffer ones, giving a larger radius for the same mass. For both EOS modes, the mass and radius credible intervals are again narrower than those obtained from the EOS-agnostic approach.
For both the EOS-informed and EOS-agnostic case, the inferred geometry consists of a ring surrounding the north pole—the primary hot spot with an omission region in the middle—and a two-temperature spot in the southern hemisphere, the secondary hot spot, consisting of a superseding and ceding region (see Fig. 6 for the inferred maximum-likelihood geometry). Notably, both the CS- and PP-informed prior runs yield geometric solutions that differ in hot spot size, temperature, and location from Choudhury et al. (2024), as seen in Fig. 13. Most credible intervals for the geometry parameters—especially those of the secondary hot spot—do not overlap, comparing the EOS-agnostic with the EOS-informed cases. The two EOS-informed runs are dominated by modes that are more extreme than the EOS-agnostic configuration, in terms of their surface emission patterns. In particular, the secondary superseding hot spot appears smaller, hotter, and closer to the south pole, while the ceding component is relatively larger, compared to both the superseding region and the EOS-agnostic solution. The temperature of the ceding component is similar for all cases. On average, the ring is comparable in size to the CS-informed run but becomes larger and thinner (with a larger omission region) in the PP-informed run, relative to the EOS-agnostic solution. For the maximum-likelihood vector, the rings in both the CS- and PP-informed run are smaller, slightly thicker with a smaller central omission region compared to the EOS-agnostic solution.
Figure 5: Radius, compactness, and mass posterior distributions using the PSR J0437-4715 NICER dataset conditional on the CST+PDT model. Three posterior distributions are shown: the headline result from Choudhury et al. (2024) using an EOS-agnostic approach (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the PP-informed MMRR prior (blue). The EOS-agnostic analysis used MultiNest with $2\times 10^{4}$ live points and a sampling efficiency of 0.3, while the EOS-informed runs used MultiNest with $4\times 10^{3}$ live points and a sampling efficiency of 0.1. See Fig. 4 for more details about the figure elements.
Parameter
Choudhury et al.
2024 (EOS agnostic)
CS-informed
$M$-$R$ prior
$P$-informed
$M$-$R$ prior
$R_{\text{eq}}[\text{km}]$ 1.62 0.74 0.58
$M/R_{\text{eq}}$ 1.63 0.85 0.57
$MM_{\odot}$]] 0.08 0.07 0.15
Table 3: KL divergence (in bits) between the posterior and corresponding prior for each parameter, computed for the three different inferences for PSR J0437-4715 shown in Fig. 5.
Figure 6: Projections of the inferred maximum-likelihood geometries are shown for the EOS-agnostic run of Choudhury et al. (2024) (top) and for the CS-informed MRM–R prior run (bottom) using the PSR J0437–4715 NICER dataset, both conditional on the CST+PDT model. The projections are shown from the star’s equatorial plane (left) and from the observer inclination (right). The colorbar shows the normalized hot spot temperature, ranging between 0 (coolest, orange) and 1 (hottest, red). In both cases, the geometries consist of a ring surrounding the north pole and a two-temperature spot in the southern hemisphere. Note that in the CS-informed run (bottom), the small (superseding) hot spot in the southern hemisphere lies on the edge of the larger, cooler (ceding) region, whereas in the EOS-agnostic run (top) it is centered within the cooler region.
Figure 7: Radius and spot geometry posterior distributions using the PSR J0437-4715 NICER dataset conditional on the CST+PDT model. Additional details of the parameter descriptions and the prior probability distribution functions can be found in Table A1 of Choudhury et al. (2024). [^13] The 68% credible interval lower limit of the secondary superseding angular radius is -0.0023, not -0.023 as stated in that table. See Fig. 5 for more details about the figure elements.
Relative to the EOS-agnostic approach, the maximum likelihood and evidence from the CS-informed prior runs are approximately +38.8+38.8 and +39.5+39.5 units higher, respectively, while those from the PP-informed prior runs are approximately +37.0+37.0 and +33.0+33.0 units higher, measured in natural logarithm. These differences in evidence correspond to a Bayes factor of $2\ln\text{BF}\approx 79.0$ (CS) and 66.0 (PP), constituting very strong evidence in favor of the EOS-informed models, following the interpretation of Kass and Raftery (1995). Furthermore, the CS-informed run is higher than the PP-informed run by +1.7+1.7 (maximum likelihood) and +6.5+6.5 (evidence), yielding $2\ln\text{BF}\approx 13.0$ which constitutes very strong evidence for the CS-informed model.
Regarding computational cost, the EOS-agnostic prior run required approximately 783k core hours, compared to 87k and 80k core hours for the runs using the CS-informed and PP-informed priors, respectively. Note that these runs are done with different sampler settings, which makes comparing the computational time difficult.
IV.2 EOS Inference
We now examine the influence of using the EOS-informed PPM results, compared to the EOS-agnostic PPM results, for PSR J0740+6620 and PSR J0437-4715, as presented in Sec. IV.1, on the subsequent EOS inference for the PP and CS models. For the EOS inference only these PPM results are included and no additional astrophysics constraints, such as those from gravitational wave observations, are used.
Figure 8 shows the pressure-energy density and MMRR posteriors for both the PP and CS models, obtained with the EOS-agnostic or the EOS-informed PPM results. For the PP model, the EOS-informed PPM results shift the pressure-energy density posterior toward stiffer EOSs, accommodating both the tighter radius constraint for PSR J0740+6620 and the larger radii inferred for PSR J0437-4715. Compared to the EOS-agnostic PPM results, the corresponding MMRR posteriors yield markedly narrower 68% and 95% credible regions, with an overall preference for larger stellar radii. On the other hand, the CS model with EOS-informed PPM results predicts relatively lower pressures at intermediate densities and correspondingly smaller radii, thereby favoring softer EOSs. Although the EOS-informed PPM results of the PP and CS models prefer stiffer and softer EOSs, respectively, both model posteriors exhibit some overlap. For instance, in the $M$-$R$ space, both EOS-informed PPM results overlap between 12 km and 12.6 km. However, when all NICER sources and gravitational-wave constraints are included, the PP and CS posteriors overlap strongly, demonstrating the overall consistency between the two models Rutherford et al. (2024).
[…]
Figure 9 shows the posterior pressure distributions at $2,3,\,\text{and}\,4n_{0}$, illustrating the impact of using EOS-informed PPM results compared to EOS-agnostic PPM results on intermediate-density EOS inferences. For the CS model, the pressures across all three fixed densities shift towards lower central values. We also find that at $3n_{0}$ and $4n_{0}$, the central pressures are more peaked than their EOS-agnostic counterpart. On the other hand, for the PP model the pressure shifts towards higher values across all three densities, with narrower and more sharply peaked distributions at $3n_{0}$ and $4n_{0}$ when including EOS-informed PPM results compared to EOS-agnostic PPM results in the EOS inference. Similar to the results in Fig. 8, the EOS-informed PPM pressure posteriors show some overlap across all three fixed densities between the PP and CS models.
Figure 9: Posterior (colored shaded regions) and prior (grey) pressure distributions at $2n_{0}$ (upper panels), $3n_{0}$ (middle panels), and $4n_{0}$ (lower panels). Results are shown for the PP (left) and CS (right) models using the EOS-agnostic (light blue) and EOS-informed (dark blue) PPM results.
V Discussion and Future Work
[…]
V.1 PPM including EOS Prior Information
Our main goal was to implement the EOS-informed MMRR priors within the PPM framework and, in doing so, obtain tighter constraints on neutron star (geometry) parameters while reducing computational cost. This objective was successfully met: for both pulsars, the credible intervals derived with EOS-informed priors are narrower than those from the EOS-agnostic approach, and the posteriors behave as expected, with the CS (PP) model favoring somewhat smaller (larger) radii relative to the EOS-agnostic case. As anticipated, the computational costs for PSR J0740+6620 were significantly reduced when using EOS-informed priors, due to the smaller parameter space the sampler had to explore. For PSR J0437-4715, however, a direct comparison of computational costs cannot be made because different sampler settings were used in the EOS-agnostic and EOS-informed runs in order to reduce runtime.
V.1.1 PSR J0740+6620
For PSR J0740+6620, the inferred masses for all three models are almost identical, with the mass posteriors being almost completely dominated by the mass prior from radio timing and the corresponding KL divergences close to zero. This outcome is as expected, given the mass prior from radio timing is highly informative.
Regarding the radius, the credible intervals overlap for all three models. Nevertheless, the median radius obtained with the PP-informed MMRR prior resembles that from the EOS-agnostic run of Salmi et al. (2024a) slightly better than the median radius from the CS-informed run. This is expected, since the PP radius prior overlaps more strongly with the EOS-agnostic posterior than the CS prior does, as seen in Fig. 4.
[…]
Apart from these differences, no particular model is clearly favored: the maximum likelihood and Bayesian evidence values for all three models are nearly identical across runs.
V.1.2 PSR J0437-4715
For PSR J0437-4715, the inferred masses for all three models are again nearly identical, with the mass posteriors largely prior-dominated and corresponding KL divergences around 0.1. Similar to PSR J0740+6620, this is expected given the highly informative mass prior from radio timing.
Concerning the radius, the credible interval obtained with the CS-informed prior overlaps with the EOS-agnostic result from Choudhury et al. (2024) at the 68% level, while the PP-informed prior only overlaps at the 95% level, as shown in Fig. 5. In both EOS-informed prior runs, the radius posterior is concentrated around the peak of the corresponding prior. However, for both the radius and the compactness posteriors, the data are clearly informative with KL divergences between 0.57 and 0.85, although the prior still contributes significantly. The EOS-agnostic model yields KL divergences of about 1.6 for both radius and compactness, indicating a substantial information gain from the data.
It is notable, however, that the CS- and PP-informed MRM–R prior runs yield significantly different geometric solutions than the one obtained in the EOS-agnostic run of Choudhury et al. (2024). In particular, the secondary dual temperature hot region is closer to the south pole in the EOS-informed runs, with the superseding hot spot smaller and hotter than in the EOS-agnostic case. This difference arises because the additional MRM–R constraints imposed by the EOS models during the PPM stage guide the sampler to explore this region of the parameter space more thoroughly. These solutions are sufficiently well separated in parameter space from those of the EOS-agnostic runs that we can consider them to be a separate mode.
V.1.3 Statistical Properties of PSR J0437-4715’s Geometry
Before discussing the physical properties of the newly identified geometric mode, we first discuss its statistical properties. In this regard, the maximum likelihood and evidence values from the CS- and PP-informed prior runs are significantly higher than those obtained in the EOS-agnostic analysis from Choudhury et al. (2024), with the CS-informed run yielding slightly larger values than the PP-informed run. To determine which mode is formally preferred in a statistical sense, however, requires both modes to be fully explored within the same converged run using the same prior.
To address this, we first revisited the EOS-agnostic run to check whether the new mode had been identified and explored in that analysis. Inspection of the samples from the EOS-agnostic run (from the Zenodo repository Choudhury, Devarshi and Salmi, Tuomo and Serena, Vinciguerra and Riley, Thomas and Kini, Yves and Watts, Anna L. and Dorsman, Bas and Bogdanov, Slavko and Guillot, Sebastien and Ray, Paul S. and Reardon, Daniel and Remillard, Ronald A. and Bilous, Anna and Huppenkothen, Daniela and Lattimer, James and Rutherford, Nathan and Arzoumanian, Zaven and Gendreau, Keith and Morsink, Sharon and Ho, Wynn C. G. (2024)), which did not use the multi-mode exploration setting, shows that MultiNest initially drew samples from regions corresponding to the CS- and PP-informed geometry modes. The sampler even continued to explore this part of parameter space to roughly 150 units below its maximum log-likelihood. However, as the run progressed, the sampler eventually concentrated on the geometric mode reported as the final result in Choudhury et al. (2024).
In principle, running with the multi-mode setting on should allow more thorough exploration of multiple different modes by locking in a select number of live points onto identified modes in proportion to the prior mass. In practice, however, this would increase the computational cost further still (the 20k live point run already required 783k core hours, and enabling multi-mode exploration requires even more live points to properly explore each identified mode). Carrying out such a run is therefore not possible with the available computational resources.
Given these limitations, we instead attempted runs with the more restricted (and hence more computationally tractable) CS- and PP-informed priors using MultiNest with the multi-mode exploration setting turned on. However, for the CS-informed prior using 4k live points and a sampling efficiency of 0.3, this resulted in seven distinct modes. One mode resembled the EOS-agnostic solution, while the six remaining modes were similar to each other in terms of most, but not all, parameters of the CS-informed run with the multi-mode exploration turned off.
By contrast, the PP-informed prior proved more tractable. With 8k live points and a sampling efficiency of 0.3, the multi-mode run identified two modes: one mode similar to the EOS-agnostic result, and one similar to the EOS-informed prior run, consistent within 95% credible level for all parameters, and within 68% credible level for most parameters. The maximum-likelihood and local evidence values for the latter mode, corresponding to the more extreme geometric solution from the EOS-informed prior run, are higher by approximately +35.3 and +29.6 units, respectively, compared to the mode resembling the EOS-agnostic result. This difference in evidence corresponds to a Bayes factor of $2\ln\text{BF}\approx 59.1$, constituting very strong evidence in favor of the new more extreme geometric mode found by the EOS-informed models, following the interpretation of Kass and Raftery (1995).
V.1.4 Physical Plausibility of PSR J0437-4715’s Geometry
We now turn to the physical properties of the newly identified geometric mode and assess whether it is physically plausible and consistent with multi-wavelength constraints from this source, while noting that more definitive conclusions on its feasibility will require testing against more detailed models and incorporating multi-wavelength observational constraints (see, e.g., recent modeling by Pétri et al., 2025).
Both the geometric mode obtained in the EOS-agnostic run and the more extreme mode identified in the EOS-informed prior runs feature hot spot geometries composed of combinations of ring-like and circular structures. These configurations can be consistent with the so-called “quadrudipolar” magnetic field configuration (Gralla et al., 2017; Chen et al., 2020). For instance, the hot spot observed near the northern magnetic pole could plausibly be explained by a nearly aligned dipolar rotator, [^14] Nearly aligned dipolar rotator refers to a neutron star whose magnetic and rotation axes are almost parallel, and whose magnetic field has a simple dipole configuration, implying two antipodal magnetic poles (polar caps). provided that a fit accommodating an additional smaller spot within the larger ring is viable. In such nearly aligned configurations, the expected current heating pattern includes a compact spot near the magnetic pole—where super-Goldreich-Julian currents are sustained by pair production—and an encircling ring formed by the volume return current (Philippov et al., 2015; Gralla et al., 2016).
In contrast, the very small hot (secondary) superseding spot observed in the newly identified mode at the southern pole is hard to reconcile with pulsar electrodynamics. Such narrow, localized structures are problematic, as they support only limited electric potential and are difficult to sustain through pair production processes (Chernoglazov et al., 2024). In this respect, the somewhat larger superseding spot for the geometric mode identified in Choudhury et al. (2024) is more physically reasonable.
To further assess whether the inferred hot spot geometry aligns with independent observational constraints, we next consider the viewing geometry of the neutron star as implied by radio observations. The geometrical priors used so far are based on the orbital inclination angle. For a narrow radio emission beam and a fully recycled millisecond pulsar binary, one would expect the orbital inclination angle to be close to the viewing angle between the magnetic axis and the observer’s line-of-sight (e.g., Kramer et al., 2021). However, the radio pulse of PSR J0437-4715 is very wide, with multiple connected components that are easily recognizable, spanning at least 200° in pulse longitude. With sufficient sensitivity, low-level radio emission can be detected covering roughly 330° of longitude. This can be explained either by a very wide radio beam, by a nearly aligned rotator (as in the new geometric mode)—which would appear to be inconsistent with the orbital inclination angle information—or by emission originating from a different source.
In addition, we consider the gamma-ray emission as additional multi-wavelength constraint on the inferred geometry. Recently, Kramer and Johnston (2025) argued that some radio emission of millisecond pulsars originates not from the polar cap but from beyond the light-cylinder, congruent with gamma-ray emission. PSR J0437-4715 indeed shows prominent gamma-ray emission, with significant overlap in pulse phase between the radio and gamma-ray pulse profiles. As argued by Kramer and Johnston (2025), this could explain both the wide longitude range of detected radio emission and the complex position angle swing that deviates from a rotating vector model. Solely considering the prominent gamma-ray emission detection suggests a combination of a large viewing angle and magnetic inclination angle, which would be inconsistent with an aligned rotator (see, e.g., Kramer and Johnston, 2025). Fitting a rotating vector model is not only extremely difficult, but also inadequate in principle if part of the radio emission does not originate from the polar cap.
A consistent picture is that both the viewing and magnetic inclination angles are large, the radio beam is intrinsically quite wide, and parts of the emission are likely originating from beyond the light cylinder. The geometrical information supplied by the orbital inclination angle is consistent with this image and remains consistent with the adopted priors. All in all, since the newly identified geometric mode in the EOS-informed prior run is close to a nearly aligned configuration, the less-aligned geometry obtained by the EOS-agnostic run is more plausible according to the gamma-ray and radio observations.
[…]
In conclusion, using EOS-informed priors in the PPM stage enabled the identification of a new mode, whose presence was already suggested by the EOS-agnostic run with multi-mode exploration setting off, but not fully explored. This mode, which has a more extreme surface emission pattern, is statistically favored over the EOS-agnostic solution by Choudhury et al. (2024). Among the EOS-informed prior runs, the CS-informed run is preferred over the PP-informed run according to the Bayes factor, suggesting that the data likely favours smaller radii even for the new geometric mode, consistent with the EOS-agnostic result. Nonetheless, serious questions remain about whether the new, more extreme geometric configuration is physical, in particular with regards to the very small (secondary superseding) hot spot. Future work should address this issue and consider modifying the geometry priors to better reflect what is physically reasonable.
V.1.5 Normalizing Flow as EOS prior
To model the EOS MMRR prior for the PPM step, we used a normalizing flow. We observed that the normalizing flow struggles to capture the MMRR distribution near $\approx 2.6$ $M_{\odot}$ and 13 km under the PP model, as shown in Fig. 2. This is likely due to sparse training samples at that boundary, which makes it more difficult for the neural network to learn the distribution edges. However, this is not a concern for our current analysis, since both pulsars have tightly constrained mass priors from radio timing which lie outside this region. We therefore do not sample from this part of the normalizing flow during the PPM step.
V.2 EOS Inference
The use of EOS priors in the PPM stage results in more stringent constraints on the inferred MRM–R posteriors of PSR J0740+6620 and PSR J0437-4715. Including these posteriors in the EOS inference further tightens the PP (CS) EOS posteriors. In the MRM–R plane, it produces a shift towards larger (smaller) radii and corresponding stiffening (softening) of the pressure-energy density relation for PP (CS), reflecting the expected behavior of these extensions, which relatively favor somewhat stiffer (softer) EOSs during the PPM inference. In particular, PSR J0437-4715 has the strongest impact on this effect.
Studying the impact of EOS-informed PPM results on intermediate-density EOS inferences shows that, for the CS model, the pressures across all three densities consistently shift toward lower values. In contrast, for the PP model the pressures shift to higher values across all three densities. Moreover, at $3n_{0}$ and $4n_{0}$, the pressure posteriors are more peaked than in the EOS-agnostic case. Overall, pressures increase for the PP model and decrease for the CS model at all densities, consistent with stiffening and softening of the EOS, respectively, when EOS-informed PPM data are used. It is interesting to note that in the EOS analysis with all NICER sources and including gravitational wave constraints, the prior differences are significantly smaller Rutherford et al. (2024).
V.3 Future Work
In future work, we aim to extend the analysis to sources such as PSR J1231-1411 Salmi et al. (2024b) and PSR J0030+0451 Salmi et al. (2023); Vinciguerra et al. (2024) which exhibit complex, multimodal solutions that are challenging to sample thoroughly within reasonable computational time. A key question is whether incorporating EOS-informed MMRR priors during the PPM stage can help mitigate these problems.
[…]
Acknowledgements.
We are grateful to Sasha Philippov for providing input to Sec. V.1.2 on the physical implications of the newly identified geometric mode of PSR J0437-4715. M.H. and A.L.W. acknowledge support from the NWO grant ENW-XL OCENW.XL21.XL21.038 Probing the phase diagram of Quantum Chromodynamics (PI: Watts). N.R. acknowledges support by NASA grant No.80NSSC22K0092. The work of N.R. was also supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). A.L.W. and D.C. acknowledge support from ERC Consolidator grant No. 865768 AEONS (PI: Watts). The work of M.M., I.S., A.S., and K.H. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020842). T.S. acknowledges funding by the Research Council of Finland grant No. 368807. The NWO Domain Science subsidized the use of the national computer facilities in this research. Part of the work was also carried out on the HELIOS cluster including dedicated nodes funded via the above-mentioned ERC Consolidator grant. This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. Software: Cython (Behnel et al., 2011), Daft (Fulford et al., 2025), GetDist (Lewis, 2025), GNU Scientific Library (Galassi et al., 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011), PyMultiNest (Buchner, 2016), Pyro Bingham et al. (2018), Python/C language (Oliphant, 2007), SciPy (Virtanen et al., 2020), NEoST (Raaijmakers et al., 2025), W&B (Biewald, 2020), and X-PSI (Riley et al., 2023). Facilities: NICER, XMM-Newton
Appendix A Supplementary Posterior Plot for PSR J0740+6620
Figure 10 shows the spot geometry and other parameter posterior distributions for PSR J0740+6620.
Figure 10: Spot geometry and other parameter posterior distributions using the PSR J0740+6620 NICER and XMM-Newton dataset conditional on the ST-U model. See Fig. 4 for more details about the figure elements.

### v2

@@
 Abstract
-NICER has enabled mass–radius inferences for pulsars using pulse profile modeling, providing constraints on the equation of state (EOS) of cold, dense matter.
+NICER has enabled mass–radius inferences for pulsars using pulse profile modeling (PPM), providing constraints on the equation of state (EOS) of cold, dense matter.
 To date, PPM and EOS inference have been carried out as two separate steps, with the former using EOS-agnostic priors.
@@
 Including the PPM posteriors in the subsequent EOS inference further tightens the mass-radius posteriors through the chiral effective field theory constraints.
-However, there is also a sensitivity to the high-density extensions, where the PP (CS) model produces a shift towards larger (smaller) radii and corresponding stiffening (softening) of the pressure–energy density relation.
+However, there is also a sensitivity to the high-density extensions, where the PP (CS) model produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure–energy density relation.
 I Introduction
@@
 (2025), [^2] https://github.com/xpsi-group/neost is then performed as a separate analysis step.
-The main motivation for this two-step approach is to provide the community with posterior samples in the MMRR plane that can be studied using any EOS model afterwards.
+The main motivation for this two-step approach is to provide the community with posterior samples in the MMRR plane that can be studied using any EOS model afterward.
 However, this strategy has some practical drawbacks:
@@
   3. 3.
-    Difficulty with multi-modality: a source like PSR J0030+0451 Salmi et al.
+    Difficulty with multimodality: A source like PSR J0030+0451 Salmi et al.
 (2023); Vinciguerra et al.
 (2024) exhibits complex, multimodal solutions that are difficult to sample thoroughly within reasonable compute time with flat priors.
-Imposing realistic priors that exclude regions already excluded by EOS constraints will focus sampling only on physically meaningful regions of parameter space, and potentially break multi-modality and reduce sampling time.
+Imposing realistic priors that exclude regions already excluded by EOS constraints will focus sampling only on physically meaningful regions of parameter space, and potentially break multimodality and reduce sampling time.
   4. 4.
@@
 In an ideal scenario, one would perform fully hierarchical Bayesian inference where the pulse profile and EOS parameters are jointly fit.
-While this would solve the aforementioned issues, the implementation of such a framework is very complex, time-consuming, and running it could be computationally intractable.
+While this would solve the aforementioned issues, the implementation of such a framework is very complex, time consuming, and running it could be computationally intractable.
 At present, the X-PSI and NEoST packages are two distinct frameworks, each focusing on different aspects of the full Bayesian model, namely PPM and EOS inference, respectively.
@@
 (2024).
-At nuclear densities beyond the BPS crust (Baym et al., 1971), these are based on next-to-next-to-next-to-leading-order (N3LO) chiral effective theory calculations of matter in beta equilibrium from Keller et al.
+At nuclear densities beyond the Baym-Pethick-Sutherland (BPS) crust (Baym et al., 1971), these are based on next-to-next-to-next-to-leading-order (N3LO) chiral effective theory calculations of matter in beta equilibrium from Keller et al.
 (2023) up to $1.5n_{0}$, where $n_{0}=0.16\,\mathrm{fm}^{-3}$ is the nuclear saturation density.
-The EOS-prior thus builds in knowledge of the crust and of the EOS at nuclear densities based on our modern understanding of nuclear forces.
+The EOS prior thus builds in knowledge of the crust and of the EOS at nuclear densities based on our modern understanding of nuclear forces.
 At higher densities, we use two possible model extensions, based on the speed of sound inside the neutron star interior (CS) Greif et al.
@@
 Sections II.3 and II.4 detail the X-PSI PPM analysis pipeline and the implementation of normalizing flows to represent the MRM–R prior derived from the two EOS models, respectively.
-Complete information of each PPM and EOS inference run, including the exact X-PSI and NEoST versions, data products, posterior sample files, and all the analysis files are available in a Zenodo repository.
-[^3] https://doi.org/10.5281/zenodo.17257403
+Complete information of each PPM and EOS inference run, including the exact X-PSI and NEoST versions, data products, posterior sample files, and all the analysis files are available in a Zenodo repository HoogkamerRutherford2025-Zenodo.
+[^3] https://doi.org/10.5281/zenodo.17257402
 II.1 Full hierarchical bayesian model
@@
 A key advantage of this approach is the occurrence of Bayesian shrinkage: By constraining all neutron stars to lie on a common EOS curve, we should obtain tighter constraints on the MRM–R relation and the EOS parameters.
-We begin by representing the pulse profile data as $\left\{D_{k,j}\right\}^{K,J}_{k,j=1}$, where $D_{k,j}$ denotes the observed data for the $k$-th neutron star with $K$ the total number of neutron stars included in the analysis, and the $j$-th data point with $J$ the total number of data points in each pulse profile.
-See Tab.
-1 for an overview of all parameter definitions.
+We begin by representing the pulse profile data as $\left\{D_{k,j}\right\}^{K,J}_{k,j=1}$, where $D_{k,j}$ denotes the observed data for the $k$th neutron star with $K$ the total number of neutron stars included in the analysis, and the $j$th data point with $J$ the total number of data points in each pulse profile.
+See Table 1 for an overview of all parameter definitions.
 Figure 1: Probabilistic graphical model illustrating the hierarchical Bayesian framework for the joint inference of neutron star properties and EOS parameters.
-Arrows denote statistical dependencies between variables.
-The shaded gray circle represent observed variables, while double circles indicate deterministic variables.
+Arrows denote statistical dependences between variables.
+The shaded gray circle represents observed variables, while double circles indicate deterministic variables.
 The box represents a summation that applies to all parameters contained within it.
-See Tab.
-1 for a detailed description of all variables.
+See Table 1 for a detailed description of all variables.
 Next, we specify the prior distribution for the parameters not directly affected by the EOS, such as the inclination angle and the hot spot geometry parameters:
@@
 In contrast, the neutron star parameters that are affected by the EOS are governed by a deterministic relationship.
-Using ($\varphi,\varepsilon_{c,k}$), the vector containing the EOS parameters and the central density of the $k$-th neutron star, respectively, and solving the Tolman-Oppenheimer-Volkoff (TOV; see Tolman, 1939; Oppenheimer and Volkoff, 1939) equations, the radius $R_{k}$ is determined by the MMRR relationship of the EOS for a given sample from the prior for $M_{k}$.
-$M_{k}$ is sampled from a mass prior based on radio timing observations (when available), or from a uniform prior (when such observations are unavailable), and $R_{k}$ is determined using $\varphi$, $\varepsilon_{c,k}$ and $M_{k}$.
+Using ($\varphi,\varepsilon_{c,k}$), the vector containing the EOS parameters and the central density of the $k$th neutron star, respectively, and solving the Tolman-Oppenheimer-Volkoff (TOV; see Tolman, 1939; Oppenheimer and Volkoff, 1939) equations, the radius $R_{k}$ is determined by the MMRR relationship of the EOS for a given sample from the prior for $M_{k}$.
+$M_{k}$ is sampled from a mass prior based on radio-timing observations (when available), or from a uniform prior (when such observations are unavailable), and $R_{k}$ is determined using $\varphi$, $\varepsilon_{c,k}$ and $M_{k}$.
 In this setup, the mass $M_{k}$ is treated as an independent variable, and we can express its prior as
@@
 $p(\varphi)\,.$ (3)
-Similarly, the prior on the central energy density $\varepsilon_{c,k}$ is given by:
+Similarly, the prior on the central energy density $\varepsilon_{c,k}$ is given by
 $p(\left\{\varepsilon_{c,k}\right\}^{K}_{k=1})=\prod_{k=1}^{K}p(\varepsilon_{c,k})\,$, (4)
@@
 […]
-where we assume that both neutron stars and data points within each pulse profile are statistically independent.
-By combining Eqs.
-(1), (2), (3), (4), and (5), we obtain the posterior probability distribution:
+where we assume that both neutron stars and data points within each pulse profile are statistically independent. By combining Eqs. (1)-(5), we obtain the posterior probability distribution:
 $p(\left\{\theta_{k},M_{k},\varepsilon_{c,k}\right\}^{K}_{k=1},\varphi|\left\{D_{k,j}\right\}^{K,J}_{k,j=1})$
@@
 $\times~p(\theta_{k})p(M_{k})p(\varepsilon_{c,k})p(\varphi)\,.$ (6)
-As mentioned before, since the equatorial radius $R_{k}$ of a given source is completely determined by the combination of $(\varphi,\varepsilon_{c,k},M_{k})$, the posterior distribution on $R_{k}$ can be computed once the posteriors on $\varphi$, $\varepsilon_{c,k}$, and $M_{k}$, are determined.
+As mentioned before, since the equatorial radius $R_{k}$ of a given source is completely determined by the combination of $(\varphi,\varepsilon_{c,k},M_{k})$, the posterior distribution on $R_{k}$ can be computed once the posteriors on $\varphi$, $\varepsilon_{c,k}$, and $M_{k}$ are determined.
 II.2 Intermediate step: PPM including EOS prior information and EOS inferencing
@@
 (2025).
-[^5] The pre-release form of NEoST was used in (Greif et al., 2019; Raaijmakers et al., 2019, 2020, 2021; Rutherford et al., 2023).
+[^5] The prerelease form of NEoST was used in (Greif et al., 2019; Raaijmakers et al., 2019, 2020, 2021; Rutherford et al., 2023).
 The released version of NEoST has been used in Rutherford et al.
@@
 (2024).
-The two-step process implemented here naturally raises concerns about double-counting the EOS prior, since it is used both to define priors for MMRR during the PPM stage and again in the subsequent EOS inference.
-This is especially relevant because uniform priors on the EOS parameters lead to non-uniform priors on neutron star mass and radius due to the non-linearity of the TOV equations.
+The two-step process implemented here naturally raises concerns about double counting the EOS prior, since it is used both to define priors for MMRR during the PPM stage and again in the subsequent EOS inference.
+This is especially relevant because uniform priors on the EOS parameters lead to nonuniform priors on neutron star mass and radius due to the nonlinearity of the TOV equations.
 However, since these equations are deterministic, the transition from EOS parameters to MMRR space corresponds to a change of variables, which is a common operation in Bayesian inference especially when it reduces the computational costs of sampling.
@@
 In this work, $\bm{d}$ consists of the MRM–R PPM posteriors for PSR J0740+6620 and PSR J0437-4715, obtained using an EOS-informed MRM–R prior together with a mass prior derived from radio timing.
-We do not include other data, such as gravitational wave observations, in the EOS inference.
+We do not include other data, such as gravitational-wave observations, in the EOS inference.
 It is important to note that when an EOS-informed prior is adopted in the PPM step, the EOS inference is only self-consistent if the same EOS model is used in both PPM and EOS inference.
-EOS inference with a different EOS model family than that used to define MMRR priors for PPM is still possible, provided that the EOS model used to define the PPM priors encompasses the full parameter range of the EOS model applied during the EOS inference.
+EOS inference with a different EOS model family from that used to define MMRR priors for PPM is still possible, provided that the EOS model used to define the PPM priors encompasses the full parameter range of the EOS model applied during the EOS inference.
 In this scenario, EOS inference is akin to using PPM results derived using EOS-agnostic priors.
@@
 If the priors for the EOS model parameters are flat, similar to what is done in this work, then the EOS inference will not be affected by the shape of those priors.
-However, if the EOS priors are non-uniform, then there is an effective double counting of the prior contribution of the EOS model, thus making the EOS posterior inference statistically inconsistent.
+However, if the EOS priors are nonuniform, then there is an effective double counting of the prior contribution of the EOS model, thus making the EOS posterior inference statistically inconsistent.
 This ensures that the prior informing the PPM-derived MRM–R distributions—the EOS parameters $\varphi$—remains aligned with the model used in the subsequent EOS inference.
@@
 (2009, 2019) and its Python binding PyMultiNest Buchner (2016) to compute posterior samples and calculate the evidence for Bayesian model comparison.
-We opted for this set-up over UltraNest Buchner (2021) due to its lower computational cost, despite UltraNest’s proven convergence and stable results for PSR J0740+6620 Hoogkamer et al.
+We opted for this setup over UltraNest Buchner (2021) due to its lower computational cost, despite UltraNest’s proven convergence and stable results for PSR J0740+6620 Hoogkamer et al.
 (2025).
@@
 (2024) (see Sec.
-II.4 for more details), as well as NICER background models to constrain the non-source background, cross-validated with XMM-Newton data.
+II.4 for more details), as well as NICER background models to constrain the nonsource background, cross-validated with XMM-Newton data.
 Instead of 20k live points and a sampling efficiency of 0.3 used by Choudhury et al.
@@
 II.4 Normalizing flow as EOS prior
-Uniform sampling in the EOS space induces a non-uniform prior in the MMRR space due to the non-linearity of the TOV equations.
+Uniform sampling in the EOS space induces a nonuniform prior in the MMRR space due to the nonlinearity of the TOV equations.
 Because an analytical transformation of the probability density between these spaces is intractable, we approximate the MMRR prior to enable sampling from it during the PPM stage.
@@
 Normalizing flows approximate the target distribution by transforming a simple and analytically known distribution, such as a standard normal distribution, into a more complex one by applying a sequence of invertible (i.e., bijective) and differentiable transformations.
-These transformations are parameterized by neural networks to allow for greater flexibility, and enable both sampling and evaluation of the log-density of the resulting distribution.
+These transformations are parametrized by neural networks to allow for greater flexibility, and enable both sampling and evaluation of the log-density of the resulting distribution.
 In this work, we use normalizing flows to model the EOS MMRR prior distribution,
@@
 allowing us to efficiently sample joint MMRR combinations, as well as draw radius samples conditioned on a given mass.
-The latter is particularly useful in cases where a mass prior is available from radio timing observations, as for PSR J0740+6620 and PSR J0437-4715.
+The latter is particularly useful in cases where a mass prior is available from radio-timing observations, as for PSR J0740+6620 and PSR J0437-4715.
 In this paper, we use the implementation of normalizing flows provided in Pyro Bingham et al.
@@
 The resulting network contains two hidden layers with 32 neurons each, and ReLU [^10] ReLU (rectified linear unit) is an activation function defined as $f(x)=\text{max}(0,x)$.
-Activation functions introduce non-linearity into the network, allowing the model to capture complex patterns and relationships in the data.
+Activation functions introduce nonlinearity into the network, allowing the model to capture complex patterns and relationships in the data.
 Nair and Hinton (2010) activations following each layer.
@@
 To evaluate, we compare the probability distribution learned by the normalizing flow to our test set using the Kullback-Leibler (KL) divergence, which provides a single number quantifying how closely the entire learned distribution approximates the target one.
-This results in KL-divergences of 0.0075 and 0.0023 for the normalizing flow of the CS and PP EOS prior distribution, respectively.
-A KL-divergence close to zero indicates that the distribution learned by the normalizing flow is nearly indistinguishable from the test set.
+This results in KL divergences of 0.0075 and 0.0023 for the normalizing flow of the CS and PP EOS prior distribution, respectively.
+A KL divergence close to zero indicates that the distribution learned by the normalizing flow is nearly indistinguishable from the test set.
 See Fig.
@@
 The bounds are conservative and lie well outside the support of the EOS prior.
-For a sample size of one million, the probability of drawing samples outside these bounds is close to zero, so truncation is not expected to significantly affect the normalization of the distribution.
+For a sample size of $1\times 10^{6}$, the probability of drawing samples outside these bounds is close to zero, so truncation is not expected to significantly affect the normalization of the distribution.
 III Data
@@
 III.1, we describe the most important aspects of the dataset from the high-mass pulsar PSR J0740+6620, as detailed in Salmi et al.
-(2024a); Hoogkamer et al.
+(2024a), and Hoogkamer et al.
 (2025).
@@
 This dataset owes its name to the usage of the 3C50 background model, which is designed to estimate the instrumental background in NICER observations.
-The model utilizes NICER event data to predict the background level as a function of energy (see Remillard et al.
+The model utilizes NICER event data to predict the background level as a function of the energy (see Remillard et al.
 (2022) for the model definition).
@@
 (2022) for a detailed description of the analysis and filtering procedure).
-The filtering steps reduce the usable exposure time from 2.736 Ms to 1.328 Ms, but enable the quantification of the uncertainties in the net background spectrum, which is predominantly influenced by systematic instrumental effects.
-To maximize the signal-to-noise ratio and mitigate contamination from the bright Seyfert II active galactic nucleus RX J0437-4711, located $4^{\prime}.18$ from PSR J0437-4715, NICER observations were conducted pointing $1^{\prime}.15$ south-west of the pulsar.
+The filtering steps reduce the usable exposure time from 2.736 to 1.328 Ms, but enable the quantification of the uncertainties in the net background spectrum, which is predominantly influenced by systematic instrumental effects.
+To maximize the signal-to-noise ratio and mitigate contamination from the bright Seyfert II active galactic nucleus RX J0437-4711, located $4^{\prime}.18$ from PSR J0437-4715, NICER observations were conducted pointing $1^{\prime}.15$ southwest of the pulsar.
 This deliberate offset reduces the AGN flux within the field of view to only $\sim$2.25% (i.e., $\sim 0.2$ counts/s).
@@
 The priors on these parameters are given by: $\Gamma_{1}\in[0,8]$, $\Gamma_{2}\in[0,8]$ $\Gamma_{3}\in[0.5,8]$, and $2n_{0}\leq n_{11}\leq n_{23}\leq 8.3n_{0}$.
-For the CS model, the prior distribution is defined by parameterizing the speed of sound $c_{s}^{2}=dP/d\varepsilon$, using five parameters $a_{1}$ to $a_{5}$, which are varied over the following ranges: $a_{1}\in[0.1,1.5]$, $a_{2}\in[1.5,12]$, $a_{3}/a_{2}\in[0.05,2]$, $a_{4}\in[1.5,37]$, and $a_{5}\in[0.1,1]$.
+For the CS model, the prior distribution is defined by parametrizing the speed of sound $c_{s}^{2}=dP/d\varepsilon$, using five parameters $a_{1}$ to $a_{5}$, which are varied over the following ranges: $a_{1}\in[0.1,1.5]$, $a_{2}\in[1.5,12]$, $a_{3}/a_{2}\in[0.05,2]$, $a_{4}\in[1.5,37]$, and $a_{5}\in[0.1,1]$.
 The resulting prior distributions for both the PP and CS models span a large range of possible EOSs.
@@
 Figure 4 shows the posterior distributions, as well as the ones obtained when incorporating the CS and PP models as prior information on MRM–R.
-Tab.
-2 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior for each model.
+Table 2 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior for each model.
 Figure 4: Radius, compactness, and mass posterior distributions using the PSR J0740+6620 joint NICER and XMM-Newton dataset conditional on the ST-U model.
@@
 10 of the Appendix, match even better for both the EOS-agnostic approach and the EOS-informed prior runs, with overlapping credible intervals for all parameters.
-Furthermore, the difference between the maximum likelihood and Bayesian evidence between all three models is negligible.
+Furthermore, the difference between the maximum likelihood and Bayesian evidence among all three models is negligible.
 [^11] The difference in maximum log-likelihood is approximately $0.0$ across all three cases, while the evidence differs by approximately +0.30+0.30 for the CS-informed run and +0.36+0.36 for the PP-informed run relative to the EOS-agnostic approach.
@@
 5.
-Tab.
-3 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior.
+Table 3 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior.
 As shown in Fig.
@@
 For both EOS modes, the mass and radius credible intervals are again narrower than those obtained from the EOS-agnostic approach.
-For both the EOS-informed and EOS-agnostic case, the inferred geometry consists of a ring surrounding the north pole—the primary hot spot with an omission region in the middle—and a two-temperature spot in the southern hemisphere, the secondary hot spot, consisting of a superseding and ceding region (see Fig.
-6 for the inferred maximum-likelihood geometry).
+For both the EOS-informed and EOS-agnostic cases, the inferred geometry consists of a ring surrounding the north pole—the primary hot spot with an omission region in the middle—and a two-temperature spot in the southern hemisphere, the secondary hot spot, consisting of a superseding and ceding region (see Fig.
+6 for the inferred maximum likelihood geometry).
 Notably, both the CS- and PP-informed prior runs yield geometric solutions that differ in hot spot size, temperature, and location from Choudhury et al.
@@
 On average, the ring is comparable in size to the CS-informed run but becomes larger and thinner (with a larger omission region) in the PP-informed run, relative to the EOS-agnostic solution.
-For the maximum-likelihood vector, the rings in both the CS- and PP-informed run are smaller, slightly thicker with a smaller central omission region compared to the EOS-agnostic solution.
+For the maximum likelihood vector, the rings in both the CS- and PP-informed runs are smaller, slightly thicker with a smaller central omission region compared to the EOS-agnostic solution.
 Figure 5: Radius, compactness, and mass posterior distributions using the PSR J0437-4715 NICER dataset conditional on the CST+PDT model.
@@
 Table 3: KL divergence (in bits) between the posterior and corresponding prior for each parameter, computed for the three different inferences for PSR J0437-4715 shown in Fig. 5.
-Figure 6: Projections of the inferred maximum-likelihood geometries are shown for the EOS-agnostic run of Choudhury et al.
+Figure 6: Projections of the inferred maximum likelihood geometries are shown for the EOS-agnostic run of Choudhury et al.
 (2024) (top) and for the CS-informed MRM–R prior run (bottom) using the PSR J0437–4715 NICER dataset, both conditional on the CST+PDT model.
@@
 IV.1, on the subsequent EOS inference for the PP and CS models.
-For the EOS inference only these PPM results are included and no additional astrophysics constraints, such as those from gravitational wave observations, are used.
-Figure 8 shows the pressure-energy density and MMRR posteriors for both the PP and CS models, obtained with the EOS-agnostic or the EOS-informed PPM results.
+For the EOS inference only these PPM results are included and no additional astrophysics constraints, such as those from gravitational-wave observations, are used.
+Figure 8 shows the pressure-energy density and MMRR posteriors for both the PP and CS models obtained with the EOS-agnostic or the EOS-informed PPM results.
 For the PP model, the EOS-informed PPM results shift the pressure-energy density posterior toward stiffer EOSs, accommodating both the tighter radius constraint for PSR J0740+6620 and the larger radii inferred for PSR J0437-4715.
@@
 Although the EOS-informed PPM results of the PP and CS models prefer stiffer and softer EOSs, respectively, both model posteriors exhibit some overlap.
-For instance, in the $M$-$R$ space, both EOS-informed PPM results overlap between 12 km and 12.6 km.
-However, when all NICER sources and gravitational-wave constraints are included, the PP and CS posteriors overlap strongly, demonstrating the overall consistency between the two models Rutherford et al.
+For instance, in the $M$-$R$ space, both EOS-informed PPM results overlap between 12 and 12.6 km.
+However, when all NICER sources and gravitational-wave constraints are included, the PP and CS posteriors overlap strongly, demonstrating the overall consistency between the two models by Rutherford et al.
 (2024).
@@
 Figure 9 shows the posterior pressure distributions at $2,3,\,\text{and}\,4n_{0}$, illustrating the impact of using EOS-informed PPM results compared to EOS-agnostic PPM results on intermediate-density EOS inferences.
-For the CS model, the pressures across all three fixed densities shift towards lower central values.
+For the CS model, the pressures across all three fixed densities shift toward lower central values.
 We also find that at $3n_{0}$ and $4n_{0}$, the central pressures are more peaked than their EOS-agnostic counterpart.
-On the other hand, for the PP model the pressure shifts towards higher values across all three densities, with narrower and more sharply peaked distributions at $3n_{0}$ and $4n_{0}$ when including EOS-informed PPM results compared to EOS-agnostic PPM results in the EOS inference.
+On the other hand, for the PP model the pressure shifts toward higher values across all three densities, with narrower and more sharply peaked distributions at $3n_{0}$ and $4n_{0}$ when including EOS-informed PPM results compared to EOS-agnostic PPM results in the EOS inference.
 Similar to the results in Fig.
 8, the EOS-informed PPM pressure posteriors show some overlap across all three fixed densities between the PP and CS models.
-Figure 9: Posterior (colored shaded regions) and prior (grey) pressure distributions at $2n_{0}$ (upper panels), $3n_{0}$ (middle panels), and $4n_{0}$ (lower panels).
+Figure 9: Posterior (colored shaded regions) and prior (gray) pressure distributions at $2n_{0}$ (upper panels), $3n_{0}$ (middle panels), and $4n_{0}$ (lower panels).
 Results are shown for the PP (left) and CS (right) models using the EOS-agnostic (light blue) and EOS-informed (dark blue) PPM results.
@@
 As anticipated, the computational costs for PSR J0740+6620 were significantly reduced when using EOS-informed priors, due to the smaller parameter space the sampler had to explore.
-For PSR J0437-4715, however, a direct comparison of computational costs cannot be made because different sampler settings were used in the EOS-agnostic and EOS-informed runs in order to reduce runtime.
+For PSR J0437-4715, however, a direct comparison of computational costs could not be made because different sampler settings were used in the EOS-agnostic and EOS-informed runs in order to reduce run-time.
 V.1.1 PSR J0740+6620
@@
 V.1.2 PSR J0437-4715
-For PSR J0437-4715, the inferred masses for all three models are again nearly identical, with the mass posteriors largely prior-dominated and corresponding KL divergences around 0.1.
+For PSR J0437-4715, the inferred masses for all three models are again nearly identical, with the mass posteriors largely prior dominated and corresponding KL divergences around 0.1.
 Similar to PSR J0740+6620, this is expected given the highly informative mass prior from radio timing.
@@
 The EOS-agnostic model yields KL divergences of about 1.6 for both radius and compactness, indicating a substantial information gain from the data.
-It is notable, however, that the CS- and PP-informed MRM–R prior runs yield significantly different geometric solutions than the one obtained in the EOS-agnostic run of Choudhury et al.
+It is notable, however, that the CS- and PP-informed MRM–R prior runs yield significantly different geometric solutions from the one obtained in the EOS-agnostic run of Choudhury et al.
 (2024).
@@
 G.
-(2024)), which did not use the multi-mode exploration setting, shows that MultiNest initially drew samples from regions corresponding to the CS- and PP-informed geometry modes.
+(2024)), which did not use the multimode exploration setting, shows that MultiNest initially drew samples from regions corresponding to the CS- and PP-informed geometry modes.
 The sampler even continued to explore this part of parameter space to roughly 150 units below its maximum log-likelihood.
@@
 (2024).
-In principle, running with the multi-mode setting on should allow more thorough exploration of multiple different modes by locking in a select number of live points onto identified modes in proportion to the prior mass.
-In practice, however, this would increase the computational cost further still (the 20k live point run already required 783k core hours, and enabling multi-mode exploration requires even more live points to properly explore each identified mode).
+In principle, running with the multimode setting on should allow more thorough exploration of multiple different modes by locking in a select number of live points onto identified modes in proportion to the prior mass.
+In practice, however, this would increase the computational cost further still (the 20k live point run already required 783k core hours, and enabling multimode exploration requires even more live points to properly explore each identified mode).
 Carrying out such a run is therefore not possible with the available computational resources.
-Given these limitations, we instead attempted runs with the more restricted (and hence more computationally tractable) CS- and PP-informed priors using MultiNest with the multi-mode exploration setting turned on.
+Given these limitations, we instead attempted runs with the more restricted (and hence more computationally tractable) CS- and PP-informed priors using MultiNest with the multimode exploration setting turned on.
 However, for the CS-informed prior using 4k live points and a sampling efficiency of 0.3, this resulted in seven distinct modes.
-One mode resembled the EOS-agnostic solution, while the six remaining modes were similar to each other in terms of most, but not all, parameters of the CS-informed run with the multi-mode exploration turned off.
+One mode resembled the EOS-agnostic solution, while the six remaining modes were similar to each other in terms of most, but not all, parameters of the CS-informed run with the multimode exploration turned off.
 By contrast, the PP-informed prior proved more tractable.
-With 8k live points and a sampling efficiency of 0.3, the multi-mode run identified two modes: one mode similar to the EOS-agnostic result, and one similar to the EOS-informed prior run, consistent within 95% credible level for all parameters, and within 68% credible level for most parameters.
-The maximum-likelihood and local evidence values for the latter mode, corresponding to the more extreme geometric solution from the EOS-informed prior run, are higher by approximately +35.3 and +29.6 units, respectively, compared to the mode resembling the EOS-agnostic result.
+With 8k live points and a sampling efficiency of 0.3, the multimode run identified two modes: one mode similar to the EOS-agnostic result, and one similar to the EOS-informed prior run, consistent within 95% credible level for all parameters, and within 68% credible level for most parameters.
+The maximum likelihood and local evidence values for the latter mode, corresponding to the more extreme geometric solution from the EOS-informed prior run, are higher by approximately +35.3 and +29.6 units, respectively, compared to the mode resembling the EOS-agnostic result.
 This difference in evidence corresponds to a Bayes factor of $2\ln\text{BF}\approx 59.1$, constituting very strong evidence in favor of the new more extreme geometric mode found by the EOS-informed models, following the interpretation of Kass and Raftery (1995).
 V.1.4 Physical plausibility of PSR J0437-4715’s geometry
-We now turn to the physical properties of the newly identified geometric mode and assess whether it is physically plausible and consistent with multi-wavelength constraints from this source, while noting that more definitive conclusions on its feasibility will require testing against more detailed models and incorporating multi-wavelength observational constraints (see, e.g., recent modeling by Pétri et al., 2025).
-Both the geometric mode obtained in the EOS-agnostic run and the more extreme mode identified in the EOS-informed prior runs feature hot spot geometries composed of combinations of ring-like and circular structures.
+We now turn to the physical properties of the newly identified geometric mode and assess whether it is physically plausible and consistent with multiwavelength constraints from this source, while noting that more definitive conclusions on its feasibility will require testing against more detailed models and incorporating multiwavelength observational constraints (see, e.g., recent modeling by Pétri et al., 2025).
+Both the geometric mode obtained in the EOS-agnostic run and the more extreme mode identified in the EOS-informed prior runs feature hot spot geometries composed of combinations of ringlike and circular structures.
 These configurations can be consistent with the so-called “quadrudipolar” magnetic field configuration (Gralla et al., 2017; Chen et al., 2020).
@@
 The geometrical priors used so far are based on the orbital inclination angle.
-For a narrow radio emission beam and a fully recycled millisecond pulsar binary, one would expect the orbital inclination angle to be close to the viewing angle between the magnetic axis and the observer’s line-of-sight (e.g., Kramer et al., 2021).
+For a narrow radio emission beam and a fully recycled millisecond pulsar binary, one would expect the orbital inclination angle to be close to the viewing angle between the magnetic axis and the observer’s line of sight (e.g., Kramer et al., 2021).
 However, the radio pulse of PSR J0437-4715 is very wide, with multiple connected components that are easily recognizable, spanning at least 200° in pulse longitude.
@@
 This can be explained either by a very wide radio beam, by a nearly aligned rotator (as in the new geometric mode)—which would appear to be inconsistent with the orbital inclination angle information—or by emission originating from a different source.
-In addition, we consider the gamma-ray emission as additional multi-wavelength constraint on the inferred geometry.
-Recently, Kramer and Johnston (2025) argued that some radio emission of millisecond pulsars originates not from the polar cap but from beyond the light-cylinder, congruent with gamma-ray emission.
+In addition, we consider the gamma-ray emission as additional multiwavelength constraint on the inferred geometry.
+Recently, Kramer and Johnston (2025) argued that some radio emission of millisecond pulsars originates not from the polar cap but from beyond the light cylinder, congruent with gamma-ray emission.
 PSR J0437-4715 indeed shows prominent gamma-ray emission, with significant overlap in pulse phase between the radio and gamma-ray pulse profiles.
@@
 […]
-In conclusion, using EOS-informed priors in the PPM stage enabled the identification of a new mode, whose presence was already suggested by the EOS-agnostic run with multi-mode exploration setting off, but not fully explored.
+In conclusion, using EOS-informed priors in the PPM stage enabled the identification of a new mode, whose presence was already suggested by the EOS-agnostic run with multimode exploration setting off, but not fully explored.
 This mode, which has a more extreme surface emission pattern, is statistically favored over the EOS-agnostic solution by Choudhury et al.
 (2024).
-Among the EOS-informed prior runs, the CS-informed run is preferred over the PP-informed run according to the Bayes factor, suggesting that the data likely favours smaller radii even for the new geometric mode, consistent with the EOS-agnostic result.
-Nonetheless, serious questions remain about whether the new, more extreme geometric configuration is physical, in particular with regards to the very small (secondary superseding) hot spot.
+Among the EOS-informed prior runs, the CS-informed run is preferred over the PP-informed run according to the Bayes factor, suggesting that the data likely favor smaller radii even for the new geometric mode, consistent with the EOS-agnostic result.
+Nonetheless, serious questions remain about whether the new, more extreme geometric configuration is physical, in particular with regard to the very small (secondary superseding) hot spot.
 Future work should address this issue and consider modifying the geometry priors to better reflect what is physically reasonable.
@@
 Including these posteriors in the EOS inference further tightens the PP (CS) EOS posteriors.
-In the MRM–R plane, it produces a shift towards larger (smaller) radii and corresponding stiffening (softening) of the pressure-energy density relation for PP (CS), reflecting the expected behavior of these extensions, which relatively favor somewhat stiffer (softer) EOSs during the PPM inference.
+In the MRM–R plane, it produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure-energy density relation for PP (CS), reflecting the expected behavior of these extensions, which relatively favor somewhat stiffer (softer) EOSs during the PPM inference.
 In particular, PSR J0437-4715 has the strongest impact on this effect.
@@
 Overall, pressures increase for the PP model and decrease for the CS model at all densities, consistent with stiffening and softening of the EOS, respectively, when EOS-informed PPM data are used.
-It is interesting to note that in the EOS analysis with all NICER sources and including gravitational wave constraints, the prior differences are significantly smaller Rutherford et al.
+It is interesting to note that in the EOS analysis with all NICER sources and including gravitational-wave constraints, the prior differences are significantly smaller Rutherford et al.
 (2024).
@@
 […]
+This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory.
+Software: Cython (Behnel et al., 2011), Daft (Fulford et al., 2025), GetDist (Lewis, 2025), GNU Scientific Library (Galassi et al., 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011), PyMultiNest (Buchner, 2016), Pyro Bingham et al.
+(2018), Python/C language (Oliphant, 2007), SciPy (Virtanen et al., 2020), NEoST (Raaijmakers et al., 2025), W&B (Biewald, 2020), and X-PSI (Riley et al., 2023).
+Facilities: NICER, XMM-Newton
 Acknowledgements.
@@
 Part of the work was also carried out on the HELIOS cluster including dedicated nodes funded via the above-mentioned ERC Consolidator grant.
-This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory.
-Software: Cython (Behnel et al., 2011), Daft (Fulford et al., 2025), GetDist (Lewis, 2025), GNU Scientific Library (Galassi et al., 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011), PyMultiNest (Buchner, 2016), Pyro Bingham et al.
-(2018), Python/C language (Oliphant, 2007), SciPy (Virtanen et al., 2020), NEoST (Raaijmakers et al., 2025), W&B (Biewald, 2020), and X-PSI (Riley et al., 2023).
-Facilities: NICER, XMM-Newton
+M.H.
+and N.R.
+contributed equally to this work and are the joint lead authors.
 Appendix A Supplementary posterior plot for PSR J0740+6620

### v3

@@
 (2024); Hoogkamer et al.
-(2025), PSR J1231-1411 Salmi et al.
+(2025a), PSR J1231-1411 Salmi et al.
 (2024b); Qi et al.
@@
 (2025).
-The resulting mass-radius (MMRR) constraints have been widely used to constrain the EOS of cold, dense matter (see, e.g., Rutherford et al., 2024; Koehn et al., 2024; Huang et al., 2025; Golomb et al., 2025; Li et al., 2025).
+The resulting mass-radius (MMRR) constraints have been widely used to constrain the EOS of cold, dense matter (see, e.g., Rutherford et al., 2024; Koehn et al., 2024; Huang et al., 2025; Golomb et al., 2025; Li et al., 2025b).
 The X-ray Pulse Simulation and Inference (X-PSI; Riley et al., 2023) code is an open-source software package for PPM and Bayesian statistical inference.
@@
 EOS inference, using the open-source software NEoST Raaijmakers et al.
-(2025), [^2] https://github.com/xpsi-group/neost is then performed as a separate analysis step.
+(2025a), [^2] https://github.com/xpsi-group/neost is then performed as a separate analysis step.
 The main motivation for this two-step approach is to provide the community with posterior samples in the MMRR plane that can be studied using any EOS model afterward.
@@
 Sections II.3 and II.4 detail the X-PSI PPM analysis pipeline and the implementation of normalizing flows to represent the MRM–R prior derived from the two EOS models, respectively.
-Complete information of each PPM and EOS inference run, including the exact X-PSI and NEoST versions, data products, posterior sample files, and all the analysis files are available in a Zenodo repository HoogkamerRutherford2025-Zenodo.
+Complete information of each PPM and EOS inference run, including the exact X-PSI and NEoST versions, data products, posterior sample files, and all the analysis files are available in a Zenodo repository Hoogkamer et al.
+(2025b).
 [^3] https://doi.org/10.5281/zenodo.17257402
@@
 To construct the EOS-informed posterior distribution for the derived neutron star properties, we begin by formulating the posterior obtained via X-PSI.
-In previous X-PSI analyses of rotation-powered millisecond pulsars (e.g., Riley et al., 2019; Salmi et al., 2023; Vinciguerra et al., 2024; Riley et al., 2021; Salmi et al., 2022, 2023, 2024a; Hoogkamer et al., 2025; Choudhury et al., 2024), analysis was done for a single star ($K=1$) and the EOS prior was not incorporated.
+In previous X-PSI analyses of rotation-powered millisecond pulsars (e.g., Riley et al., 2019; Salmi et al., 2023; Vinciguerra et al., 2024; Riley et al., 2021; Salmi et al., 2022, 2023, 2024a; Hoogkamer et al., 2025a; Choudhury et al., 2024), analysis was done for a single star ($K=1$) and the EOS prior was not incorporated.
 Instead, the posterior was given by:
@@
 (2021) Bayesian framework using NEoST (v2.1.0)Raaijmakers et al.
-(2025).
+(2025b).
 [^5] The prerelease form of NEoST was used in (Greif et al., 2019; Raaijmakers et al., 2019, 2020, 2021; Rutherford et al., 2023).
@@
 We opted for this setup over UltraNest Buchner (2021) due to its lower computational cost, despite UltraNest’s proven convergence and stable results for PSR J0740+6620 Hoogkamer et al.
-(2025).
+(2025a).
 For the massive pulsar PSR J0740+6620, we use the exact same setup and sampler settings as in the headline result of Salmi et al.
 (2024a) and Hoogkamer et al.
-(2025), where NICER and XMM-Newton data are jointly modeled.
+(2025a), where NICER and XMM-Newton data are jointly modeled.
 This setup includes informative priors on mass, inclination, and distance obtained from radio timing Fonseca et al.
@@
 (2024a), and Hoogkamer et al.
-(2025).
+(2025a).
 Section III.2 gives a brief description of the observations from the nearest and brightest rotation-powered millisecond pulsar PSR J0437-4715, as detailed in Choudhury et al.
@@
 (2024a); Hoogkamer et al.
-(2025).
+(2025a).
 We use the pulse-invariant channel subset [30, 150), corresponding to the nominal photon energy range [0.3, 1.5] keV.
@@
 (2022, 2024a); Hoogkamer et al.
-(2025)).
+(2025a)).
 Furthermore, the hot emitting regions of the neutron star are modeled with two independent circular uniform temperature spots using the ST-U (Single Temperature-Unshared) model (see Riley et al.
@@
 The radius credible intervals for all three models overlap.
-However, the radius inferred using the CS-informed prior is $11.90^{+0.60}_{-0.55}$ km—with the median smaller than the baseline—while the PP-informed prior yielded a slightly larger median radius of $12.52^{+0.54}_{-0.61}$ km.
+However, the radius inferred using the CS-informed prior is $11.94^{+0.65}_{-0.58}$ km—with the median smaller than the baseline—while the PP-informed prior yielded a slightly larger median radius of $12.51^{+0.56}_{-0.66}$ km.
 The results are in line with our expectation, since the CS model favors softer EOSs and the PP model favors stiffer ones.
@@
 Furthermore, the difference between the maximum likelihood and Bayesian evidence among all three models is negligible.
-[^11] The difference in maximum log-likelihood is approximately $0.0$ across all three cases, while the evidence differs by approximately +0.30+0.30 for the CS-informed run and +0.36+0.36 for the PP-informed run relative to the EOS-agnostic approach.
-In terms of computational cost, the EOS-agnostic inference required 84k core hours, compared to 41k and 36k core hours for the runs using the CS-informed and PP-informed priors, respectively.
+[^11] The difference in maximum log-likelihood is approximately $0.1$ across all three cases, while the evidence differs by approximately +0.51+0.51 for the CS-informed run and +0.24+0.24 for the PP-informed run relative to the EOS-agnostic approach.
+In terms of computational cost, the EOS-agnostic inference required 84k core hours, compared to 21k and 22k core hours for the runs using the CS-informed and PP-informed priors, respectively.
 Parameter
@@
 $M$-$R$ prior
-$R_{\text{eq}}[\text{km}]$ 0.64 0.38 0.15
-$M/R_{\text{eq}}$ 0.72 0.54 0.16
+$R_{\text{eq}}[\text{km}]$ 0.64 0.36 0.11
+$M/R_{\text{eq}}$ 0.72 0.51 0.14
 $MM_{\odot}$]] 0.01 0.03 0.01
@@
 5, the EOS-agnostic approach yielded an equatorial radius of $11.36^{+0.95}_{-0.63}$ km and a gravitational mass of $1.418^{+0.037}_{-0.037}$ $M_{\odot}$.
-The CS-informed prior run inferred a radius of $11.03^{+0.43}_{-0.41}$ km, which is slightly smaller than the baseline result, though the credible intervals otherwise strongly overlap.
-On the other hand, the PP-informed prior gives a radius of $12.92^{+0.25}_{-0.28}$ km, which is noticeably larger.
+The CS-informed prior run inferred a radius of $11.00^{+0.49}_{-0.48}$ km, which is slightly smaller than the baseline result, though the credible intervals otherwise strongly overlap.
+On the other hand, the PP-informed prior gives a radius of $12.81^{+0.30}_{-0.51}$ km, which is noticeably larger.
 Similar to PSR J0740+6620, these results align with expectations: The CS model favors softer EOSs given a smaller radius, while the PP model favors stiffer ones, giving a larger radius for the same mass.
@@
 $M$-$R$ prior
-$R_{\text{eq}}[\text{km}]$ 1.62 0.74 0.58
-$M/R_{\text{eq}}$ 1.63 0.85 0.57
-$MM_{\odot}$]] 0.08 0.07 0.15
+$R_{\text{eq}}[\text{km}]$ 1.62 0.59 0.28
+$M/R_{\text{eq}}$ 1.63 0.66 0.30
+$MM_{\odot}$]] 0.08 0.07 0.16
 Table 3: KL divergence (in bits) between the posterior and corresponding prior for each parameter, computed for the three different inferences for PSR J0437-4715 shown in Fig. 5.
@@
 5 for more details about the figure elements.
-Relative to the EOS-agnostic approach, the maximum likelihood and evidence from the CS-informed prior runs are approximately +38.8+38.8 and +39.5+39.5 units higher, respectively, while those from the PP-informed prior runs are approximately +37.0+37.0 and +33.0+33.0 units higher, measured in natural logarithm.
-These differences in evidence correspond to a Bayes factor of $2\ln\text{BF}\approx 79.0$ (CS) and 66.0 (PP), constituting very strong evidence in favor of the EOS-informed models, following the interpretation of Kass and Raftery (1995).
-Furthermore, the CS-informed run is higher than the PP-informed run by +1.7+1.7 (maximum likelihood) and +6.5+6.5 (evidence), yielding $2\ln\text{BF}\approx 13.0$ which constitutes very strong evidence for the CS-informed model.
-Regarding computational cost, the EOS-agnostic prior run required approximately 783k core hours, compared to 87k and 80k core hours for the runs using the CS-informed and PP-informed priors, respectively.
+Relative to the EOS-agnostic approach, the maximum likelihood and evidence from the CS-informed prior runs are both approximately +38.5+38.5 units higher, while those from the PP-informed prior runs are approximately +38.6+38.6 and +34.2+34.2 units higher, respectively, measured in natural logarithm.
+These differences in evidence correspond to a Bayes factor of $2\ln\text{BF}\approx 77.0$ (CS) and 68.4 (PP), constituting very strong evidence in favor of the EOS-informed models, following the interpretation of Kass and Raftery (1995).
+Furthermore, the CS-informed run is higher than the PP-informed run by +4.3+4.3 (evidence), yielding $2\ln\text{BF}\approx 8.7$ which constitutes strong evidence for the CS-informed model.
+Regarding computational cost, the EOS-agnostic prior run required approximately 783k core hours, compared to 33k and 25k core hours for the runs using the CS-informed and PP-informed priors, respectively.
 Note that these runs are done with different sampler settings, which makes comparing the computational time difficult.
@@
 For the CS model, the pressures across all three fixed densities shift toward lower central values.
-We also find that at $3n_{0}$ and $4n_{0}$, the central pressures are more peaked than their EOS-agnostic counterpart.
 On the other hand, for the PP model the pressure shifts toward higher values across all three densities, with narrower and more sharply peaked distributions at $3n_{0}$ and $4n_{0}$ when including EOS-informed PPM results compared to EOS-agnostic PPM results in the EOS inference.
@@
 In both EOS-informed prior runs, the radius posterior is concentrated around the peak of the corresponding prior.
-However, for both the radius and the compactness posteriors, the data are clearly informative with KL divergences between 0.57 and 0.85, although the prior still contributes significantly.
+However, for both the radius and the compactness posteriors, the data are clearly informative with KL divergences between 0.28 and 0.66, although the prior still contributes significantly.
 The EOS-agnostic model yields KL divergences of about 1.6 for both radius and compactness, indicating a substantial information gain from the data.
@@
 In contrast, for the PP model the pressures shift to higher values across all three densities.
-Moreover, at $3n_{0}$ and $4n_{0}$, the pressure posteriors are more peaked than in the EOS-agnostic case.
+Moreover, at $3n_{0}$ and $4n_{0}$, the PP-informed pressure posteriors are more peaked than in the EOS-agnostic case.
 Overall, pressures increase for the PP model and decrease for the CS model at all densities, consistent with stiffening and softening of the EOS, respectively, when EOS-informed PPM data are used.
@@
 This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory.
-Software: Cython (Behnel et al., 2011), Daft (Fulford et al., 2025), GetDist (Lewis, 2025), GNU Scientific Library (Galassi et al., 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011), PyMultiNest (Buchner, 2016), Pyro Bingham et al.
-(2018), Python/C language (Oliphant, 2007), SciPy (Virtanen et al., 2020), NEoST (Raaijmakers et al., 2025), W&B (Biewald, 2020), and X-PSI (Riley et al., 2023).
+Software: Cython (Behnel et al., 2011), Daft (Fulford et al., 2025), GetDist (Lewis, 2025), GNU Scientific Library (Galassi et al., 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), 2014), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011), PyMultiNest (Buchner, 2016), Pyro Bingham et al.
+(2018), Python/C language (Oliphant, 2007), SciPy (Virtanen et al., 2020), NEoST (Raaijmakers et al., 2025b), W&B (Biewald, 2020), and X-PSI (Riley et al., 2023).
 Facilities: NICER, XMM-Newton

### v5

@@
 By focusing on the parameter space consistent with certain EOSs, this approach both tightens constraints on neutron star parameters while reducing computational costs and requiring minimal additional implementation effort.
-We test this approach on two pulsars, PSR J0740+6620 and PSR J0437-4715, and with two EOS model families: a model based on the speed of sound inside the neutron star interior (CS) and a piecewise-polytropic (PP) model.
+We test this approach on two pulsars, PSR J0740+6620 and PSR J0437-4715, and with two EOS model families: a model based on the speed of sound inside the neutron star interior (CS) and a piecewise-polytropic ($P$) model.
 Both EOS models implement constraints from chiral effective field theory calculations of dense matter.
-For both pulsar datasets, the inferred radius credible intervals are narrower than in the EOS-agnostic case, with CS favoring smaller radii and PP favoring larger radii.
+For both pulsar datasets, the inferred radius credible intervals are narrower than in the EOS-agnostic case, with CS favoring smaller radii and $P$ favoring larger radii.
 For PSR J0437-4715, the EOS-informed priors reveal a new, more extreme geometric mode that is statistically favored but physically questionable.
 Including the PPM posteriors in the subsequent EOS inference further tightens the mass-radius posteriors through the chiral effective field theory constraints.
-However, there is also a sensitivity to the high-density extensions, where the PP (CS) model produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure–energy density relation.
+However, there is also a sensitivity to the high-density extensions, where the $P$ (CS) model produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure–energy density relation.
 I Introduction
@@
 At higher densities, we use two possible model extensions, based on the speed of sound inside the neutron star interior (CS) Greif et al.
-(2019) or based on a piecewise-polytropic (PP) model Hebeler et al.
+(2019) or based on a piecewise-polytropic ($P$) model Hebeler et al.
 (2013).
@@
 In Sec.
-III, we summarize the datasets for PSR J0740+6620 and PSR J0437–4715, and outline the prior intervals for the PP and CS EOS model parameters used in this analysis.
+III, we summarize the datasets for PSR J0740+6620 and PSR J0437–4715, and outline the prior intervals for the $P$ and CS EOS model parameters used in this analysis.
 Section IV presents the EOS-prior-informed PPM posteriors and the resulting constraints on the cold, dense matter EOS.
@@
 However, since these hyperparameters are fixed and therefore not random variables, we have chosen not to write them down for clarity.
-Similarly, all distributions in the joint PPM-EOS inference should be read as conditional on the choices of PPM model (see, e.g., Riley et al., 2019; Vinciguerra et al., 2023) and on the choice of EOS model (see Hebeler et al., 2013; Greif et al., 2019, for the PP and CS models, respectively).
+Similarly, all distributions in the joint PPM-EOS inference should be read as conditional on the choices of PPM model (see, e.g., Riley et al., 2019; Vinciguerra et al., 2023) and on the choice of EOS model (see Hebeler et al., 2013; Greif et al., 2019, for the $P$ and CS models, respectively).
 To avoid cluttering the notation, we have omitted including these in the conditional densities as well.
@@
 In this work, however, we focus solely on incorporating EOS prior information into the PPM-derived MRM–R posteriors, and therefore omit the gravitational-wave constraints to highlight the differences between the original and new frameworks.
-Since the EOS model parameters considered in this work—namely PP and CS—are uniformly sampled, the EOS inference framework is identical to that used in our previous work Raaijmakers et al.
+Since the EOS model parameters considered in this work—namely $P$ and CS—are uniformly sampled, the EOS inference framework is identical to that used in our previous work Raaijmakers et al.
 (2021); Rutherford et al.
@@
 Because there were many configurations with similar validation losses, we chose one with a relatively short training time.
-This configuration—using a normalizing flow with ten transformations for both $p(R|M)$ and p(M)p(M)—performed well, achieving a test loss of 2.52 for CS and 2.14 for PP (see Fig.
+This configuration—using a normalizing flow with ten transformations for both $p(R|M)$ and p(M)p(M)—performed well, achieving a test loss of 2.52 for CS and 2.14 for $P$ (see Fig.
 3 for the training and validation loss curves) in approximately 27 minutes for both cases.
@@
 The model is optimized by minimizing the negative log-likelihood of the MMRR samples.
-Figure 3: Training (orange) and validation (blue) loss curves for the PP EOS model (top) and CS EOS model (bottom).
+Figure 3: Training (orange) and validation (blue) loss curves for the $P$ EOS model (top) and CS EOS model (bottom).
 To evaluate, we compare the probability distribution learned by the normalizing flow to our test set using the Kullback-Leibler (KL) divergence, which provides a single number quantifying how closely the entire learned distribution approximates the target one.
-This results in KL divergences of 0.0075 and 0.0023 for the normalizing flow of the CS and PP EOS prior distribution, respectively.
+This results in KL divergences of 0.0075 and 0.0023 for the normalizing flow of the CS and $P$ EOS prior distribution, respectively.
 A KL divergence close to zero indicates that the distribution learned by the normalizing flow is nearly indistinguishable from the test set.
@@
 2 for a visual comparison between the prior model and the resultant normalizing flow.
-When sampling radii conditioned on a given mass, as in the cases of PSR J0740+6620 and PSR J0437-4715, samples with radii outside the interval 5.5 to 14.0 km for the PP EOS model and 6.5 to 14.5 km for the CS EOS model are excluded from the analysis.
+When sampling radii conditioned on a given mass, as in the cases of PSR J0740+6620 and PSR J0437-4715, samples with radii outside the interval 5.5 to 14.0 km for the $P$ EOS model and 6.5 to 14.5 km for the CS EOS model are excluded from the analysis.
 These bounds are chosen to avoid artifacts arising in regions where the normalizing flow assigns low but nonzero probability and to avoid drawing unphysical MMRR values.
@@
 In Sec.
-III.3, we summarize the prior distributions of the PP and CS EOS models.
+III.3, we summarize the prior distributions of the $P$ and CS EOS models.
 III.1 PSR J0740+6620
@@
 (2023, 2024)).
-III.3 PP and CS EOS prior distributions
+III.3 $P$ and CS EOS prior distributions
 With the EOS-informed PPM framework in place (Sec.
-II.2), the next step is to compute the MMRR priors for the PP and CS EOS models.
+II.2), the next step is to compute the MMRR priors for the $P$ and CS EOS models.
 This involves constructing the EOS models across the relevant density range and defining the prior ranges that specify their distributions.
@@
 […]
-For the PP model, the priors are defined by varying three polytropic indices, $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$, as well as two transition densities, $n_{12}$ and $n_{23}$, which connect the first to the second polytrope and the second to the third, respectively.
+For the $P$ model, the priors are defined by varying three polytropic indices, $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$, as well as two transition densities, $n_{12}$ and $n_{23}$, which connect the first to the second polytrope and the second to the third, respectively.
 The priors on these parameters are given by: $\Gamma_{1}\in[0,8]$, $\Gamma_{2}\in[0,8]$ $\Gamma_{3}\in[0.5,8]$, and $2n_{0}\leq n_{11}\leq n_{23}\leq 8.3n_{0}$.
 For the CS model, the prior distribution is defined by parametrizing the speed of sound $c_{s}^{2}=dP/d\varepsilon$, using five parameters $a_{1}$ to $a_{5}$, which are varied over the following ranges: $a_{1}\in[0.1,1.5]$, $a_{2}\in[1.5,12]$, $a_{3}/a_{2}\in[0.05,2]$, $a_{4}\in[1.5,37]$, and $a_{5}\in[0.1,1]$.
-The resulting prior distributions for both the PP and CS models span a large range of possible EOSs.
-For additional details on the PP and CS EOS models, see Rutherford et al.
+The resulting prior distributions for both the $P$ and CS models span a large range of possible EOSs.
+For additional details on the $P$ and CS EOS models, see Rutherford et al.
 (2024); Hebeler et al.
@@
 (2025).
-For completeness, the resulting MMRR prior distributions of both of the PP and CS models are shown in Fig.
+For completeness, the resulting MMRR prior distributions of both of the $P$ and CS models are shown in Fig.
 2.
@@
 IV.1 PPM including EOS prior information
-To assess the influence of an EOS-informed MRM–R prior on the inferred properties of PSR J0740+6620 and PSR J0437-4715, we evaluate two representative EOS models as priors: CS and PP.
+To assess the influence of an EOS-informed MRM–R prior on the inferred properties of PSR J0740+6620 and PSR J0437-4715, we evaluate two representative EOS models as priors: CS and $P$.
 These are compared against the baseline EOS-agnostic approach.
@@
 (2024a), using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01, reported an equatorial radius of $12.49^{+1.28}_{-0.88}$ km and a gravitational mass of $2.073^{+0.069}_{-0.069}$ $M_{\odot}$.
-Figure 4 shows the posterior distributions, as well as the ones obtained when incorporating the CS and PP models as prior information on MRM–R.
+Figure 4 shows the posterior distributions, as well as the ones obtained when incorporating the CS and $P$ models as prior information on MRM–R.
 Table 2 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior for each model.
@@
 Three posterior distributions are shown: the results from Salmi et al.
-(2024a) without an EOS-informed prior (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the PP-informed MMRR prior (blue).
+(2024a) without an EOS-informed prior (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the $P$-informed MMRR prior (blue).
 All inference results are obtained using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01.
@@
 The radius credible intervals for all three models overlap.
-However, the radius inferred using the CS-informed prior is $11.94^{+0.65}_{-0.58}$ km—with the median smaller than the baseline—while the PP-informed prior yielded a slightly larger median radius of $12.51^{+0.56}_{-0.66}$ km.
-The results are in line with our expectation, since the CS model favors softer EOSs and the PP model favors stiffer ones.
-However, once several sources are included, the CS and PP differences are smaller Rutherford et al.
+However, the radius inferred using the CS-informed prior is $11.94^{+0.65}_{-0.58}$ km—with the median smaller than the baseline—while the $P$-informed prior yielded a slightly larger median radius of $12.51^{+0.56}_{-0.66}$ km.
+The results are in line with our expectation, since the CS model favors softer EOSs and the $P$ model favors stiffer ones.
+However, once several sources are included, the CS and $P$ differences are smaller Rutherford et al.
 (2024).
@@
 Furthermore, the difference between the maximum likelihood and Bayesian evidence among all three models is negligible.
-[^11] The difference in maximum log-likelihood is approximately $0.1$ across all three cases, while the evidence differs by approximately +0.51+0.51 for the CS-informed run and +0.24+0.24 for the PP-informed run relative to the EOS-agnostic approach.
-In terms of computational cost, the EOS-agnostic inference required 84k core hours, compared to 21k and 22k core hours for the runs using the CS-informed and PP-informed priors, respectively.
+[^11] The difference in maximum log-likelihood is approximately $0.1$ across all three cases, while the evidence differs by approximately +0.51+0.51 for the CS-informed run and +0.24+0.24 for the $P$-informed run relative to the EOS-agnostic approach.
+In terms of computational cost, the EOS-agnostic inference required 84k core hours, compared to 21k and 22k core hours for the runs using the CS-informed and $P$-informed priors, respectively.
 Parameter
@@
 The CS-informed prior run inferred a radius of $11.00^{+0.49}_{-0.48}$ km, which is slightly smaller than the baseline result, though the credible intervals otherwise strongly overlap.
-On the other hand, the PP-informed prior gives a radius of $12.81^{+0.30}_{-0.51}$ km, which is noticeably larger.
-Similar to PSR J0740+6620, these results align with expectations: The CS model favors softer EOSs given a smaller radius, while the PP model favors stiffer ones, giving a larger radius for the same mass.
+On the other hand, the $P$-informed prior gives a radius of $12.81^{+0.30}_{-0.51}$ km, which is noticeably larger.
+Similar to PSR J0740+6620, these results align with expectations: The CS model favors softer EOSs given a smaller radius, while the $P$ model favors stiffer ones, giving a larger radius for the same mass.
 For both EOS modes, the mass and radius credible intervals are again narrower than those obtained from the EOS-agnostic approach.
@@
 6 for the inferred maximum likelihood geometry).
-Notably, both the CS- and PP-informed prior runs yield geometric solutions that differ in hot spot size, temperature, and location from Choudhury et al.
+Notably, both the CS- and $P$-informed prior runs yield geometric solutions that differ in hot spot size, temperature, and location from Choudhury et al.
 (2024), as seen in Fig.
@@
 The temperature of the ceding component is similar for all cases.
-On average, the ring is comparable in size to the CS-informed run but becomes larger and thinner (with a larger omission region) in the PP-informed run, relative to the EOS-agnostic solution.
-For the maximum likelihood vector, the rings in both the CS- and PP-informed runs are smaller, slightly thicker with a smaller central omission region compared to the EOS-agnostic solution.
+On average, the ring is comparable in size to the CS-informed run but becomes larger and thinner (with a larger omission region) in the $P$-informed run, relative to the EOS-agnostic solution.
+For the maximum likelihood vector, the rings in both the CS- and $P$-informed runs are smaller, slightly thicker with a smaller central omission region compared to the EOS-agnostic solution.
 Figure 5: Radius, compactness, and mass posterior distributions using the PSR J0437-4715 NICER dataset conditional on the CST+PDT model.
 Three posterior distributions are shown: the headline result from Choudhury et al.
-(2024) using an EOS-agnostic approach (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the PP-informed MMRR prior (blue).
+(2024) using an EOS-agnostic approach (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the $P$-informed MMRR prior (blue).
 The EOS-agnostic analysis used MultiNest with $2\times 10^{4}$ live points and a sampling efficiency of 0.3, while the EOS-informed runs used MultiNest with $4\times 10^{3}$ live points and a sampling efficiency of 0.1.
@@
 5 for more details about the figure elements.
-Relative to the EOS-agnostic approach, the maximum likelihood and evidence from the CS-informed prior runs are both approximately +38.5+38.5 units higher, while those from the PP-informed prior runs are approximately +38.6+38.6 and +34.2+34.2 units higher, respectively, measured in natural logarithm.
-These differences in evidence correspond to a Bayes factor of $2\ln\text{BF}\approx 77.0$ (CS) and 68.4 (PP), constituting very strong evidence in favor of the EOS-informed models, following the interpretation of Kass and Raftery (1995).
-Furthermore, the CS-informed run is higher than the PP-informed run by +4.3+4.3 (evidence), yielding $2\ln\text{BF}\approx 8.7$ which constitutes strong evidence for the CS-informed model.
-Regarding computational cost, the EOS-agnostic prior run required approximately 783k core hours, compared to 33k and 25k core hours for the runs using the CS-informed and PP-informed priors, respectively.
+Relative to the EOS-agnostic approach, the maximum likelihood and evidence from the CS-informed prior runs are both approximately +38.5+38.5 units higher, while those from the $P$-informed prior runs are approximately +38.6+38.6 and +34.2+34.2 units higher, respectively, measured in natural logarithm.
+These differences in evidence correspond to a Bayes factor of $2\ln\text{BF}\approx 77.0$ (CS) and 68.4 ($P$), constituting very strong evidence in favor of the EOS-informed models, following the interpretation of Kass and Raftery (1995).
+Furthermore, the CS-informed run is higher than the $P$-informed run by +4.3+4.3 (evidence), yielding $2\ln\text{BF}\approx 8.7$ which constitutes strong evidence for the CS-informed model.
+Regarding computational cost, the EOS-agnostic prior run required approximately 783k core hours, compared to 33k and 25k core hours for the runs using the CS-informed and $P$-informed priors, respectively.
 Note that these runs are done with different sampler settings, which makes comparing the computational time difficult.
@@
 We now examine the influence of using the EOS-informed PPM results, compared to the EOS-agnostic PPM results, for PSR J0740+6620 and PSR J0437-4715, as presented in Sec.
-IV.1, on the subsequent EOS inference for the PP and CS models.
+IV.1, on the subsequent EOS inference for the $P$ and CS models.
 For the EOS inference only these PPM results are included and no additional astrophysics constraints, such as those from gravitational-wave observations, are used.
-Figure 8 shows the pressure-energy density and MMRR posteriors for both the PP and CS models obtained with the EOS-agnostic or the EOS-informed PPM results.
-For the PP model, the EOS-informed PPM results shift the pressure-energy density posterior toward stiffer EOSs, accommodating both the tighter radius constraint for PSR J0740+6620 and the larger radii inferred for PSR J0437-4715.
+Figure 8 shows the pressure-energy density and MMRR posteriors for both the $P$ and CS models obtained with the EOS-agnostic or the EOS-informed PPM results.
+For the $P$ model, the EOS-informed PPM results shift the pressure-energy density posterior toward stiffer EOSs, accommodating both the tighter radius constraint for PSR J0740+6620 and the larger radii inferred for PSR J0437-4715.
 Compared to the EOS-agnostic PPM results, the corresponding MMRR posteriors yield markedly narrower 68% and 95% credible regions, with an overall preference for larger stellar radii.
 On the other hand, the CS model with EOS-informed PPM results predicts relatively lower pressures at intermediate densities and correspondingly smaller radii, thereby favoring softer EOSs.
-Although the EOS-informed PPM results of the PP and CS models prefer stiffer and softer EOSs, respectively, both model posteriors exhibit some overlap.
+Although the EOS-informed PPM results of the $P$ and CS models prefer stiffer and softer EOSs, respectively, both model posteriors exhibit some overlap.
 For instance, in the $M$-$R$ space, both EOS-informed PPM results overlap between 12 and 12.6 km.
-However, when all NICER sources and gravitational-wave constraints are included, the PP and CS posteriors overlap strongly, demonstrating the overall consistency between the two models by Rutherford et al.
+However, when all NICER sources and gravitational-wave constraints are included, the $P$ and CS posteriors overlap strongly, demonstrating the overall consistency between the two models by Rutherford et al.
 (2024).
@@
 For the CS model, the pressures across all three fixed densities shift toward lower central values.
-On the other hand, for the PP model the pressure shifts toward higher values across all three densities, with narrower and more sharply peaked distributions at $3n_{0}$ and $4n_{0}$ when including EOS-informed PPM results compared to EOS-agnostic PPM results in the EOS inference.
+On the other hand, for the $P$ model the pressure shifts toward higher values across all three densities, with narrower and more sharply peaked distributions at $3n_{0}$ and $4n_{0}$ when including EOS-informed PPM results compared to EOS-agnostic PPM results in the EOS inference.
 Similar to the results in Fig.
-8, the EOS-informed PPM pressure posteriors show some overlap across all three fixed densities between the PP and CS models.
+8, the EOS-informed PPM pressure posteriors show some overlap across all three fixed densities between the $P$ and CS models.
 Figure 9: Posterior (colored shaded regions) and prior (gray) pressure distributions at $2n_{0}$ (upper panels), $3n_{0}$ (middle panels), and $4n_{0}$ (lower panels).
-Results are shown for the PP (left) and CS (right) models using the EOS-agnostic (light blue) and EOS-informed (dark blue) PPM results.
+Results are shown for the $P$ (left) and CS (right) models using the EOS-agnostic (light blue) and EOS-informed (dark blue) PPM results.
 V Discussion and future work
@@
 Our main goal was to implement the EOS-informed MMRR priors within the PPM framework and, in doing so, obtain tighter constraints on neutron star (geometry) parameters while reducing computational cost.
-This objective was successfully met: For both pulsars, the credible intervals derived with EOS-informed priors are narrower than those from the EOS-agnostic approach, and the posteriors behave as expected, with the CS (PP) model favoring somewhat smaller (larger) radii relative to the EOS-agnostic case.
+This objective was successfully met: For both pulsars, the credible intervals derived with EOS-informed priors are narrower than those from the EOS-agnostic approach, and the posteriors behave as expected, with the CS ($P$) model favoring somewhat smaller (larger) radii relative to the EOS-agnostic case.
 As anticipated, the computational costs for PSR J0740+6620 were significantly reduced when using EOS-informed priors, due to the smaller parameter space the sampler had to explore.
@@
 Regarding the radius, the credible intervals overlap for all three models.
-Nevertheless, the median radius obtained with the PP-informed MMRR prior resembles that from the EOS-agnostic run of Salmi et al.
+Nevertheless, the median radius obtained with the $P$-informed MMRR prior resembles that from the EOS-agnostic run of Salmi et al.
 (2024a) slightly better than the median radius from the CS-informed run.
-This is expected, since the PP radius prior overlaps more strongly with the EOS-agnostic posterior than the CS prior does, as seen in Fig.
+This is expected, since the $P$ radius prior overlaps more strongly with the EOS-agnostic posterior than the CS prior does, as seen in Fig.
 4.
@@
 Concerning the radius, the credible interval obtained with the CS-informed prior overlaps with the EOS-agnostic result from Choudhury et al.
-(2024) at the 68% level, while the PP-informed prior only overlaps at the 95% level, as shown in Fig.
+(2024) at the 68% level, while the $P$-informed prior only overlaps at the 95% level, as shown in Fig.
 5.
@@
 The EOS-agnostic model yields KL divergences of about 1.6 for both radius and compactness, indicating a substantial information gain from the data.
-It is notable, however, that the CS- and PP-informed MRM–R prior runs yield significantly different geometric solutions from the one obtained in the EOS-agnostic run of Choudhury et al.
+It is notable, however, that the CS- and $P$-informed MRM–R prior runs yield significantly different geometric solutions from the one obtained in the EOS-agnostic run of Choudhury et al.
 (2024).
@@
 Before discussing the physical properties of the newly identified geometric mode, we first discuss its statistical properties.
-In this regard, the maximum likelihood and evidence values from the CS- and PP-informed prior runs are significantly higher than those obtained in the EOS-agnostic analysis from Choudhury et al.
-(2024), with the CS-informed run yielding slightly larger values than the PP-informed run.
+In this regard, the maximum likelihood and evidence values from the CS- and $P$-informed prior runs are significantly higher than those obtained in the EOS-agnostic analysis from Choudhury et al.
+(2024), with the CS-informed run yielding slightly larger values than the $P$-informed run.
 To determine which mode is formally preferred in a statistical sense, however, requires both modes to be fully explored within the same converged run using the same prior.
@@
 G.
-(2024)), which did not use the multimode exploration setting, shows that MultiNest initially drew samples from regions corresponding to the CS- and PP-informed geometry modes.
+(2024)), which did not use the multimode exploration setting, shows that MultiNest initially drew samples from regions corresponding to the CS- and $P$-informed geometry modes.
 The sampler even continued to explore this part of parameter space to roughly 150 units below its maximum log-likelihood.
@@
 Carrying out such a run is therefore not possible with the available computational resources.
-Given these limitations, we instead attempted runs with the more restricted (and hence more computationally tractable) CS- and PP-informed priors using MultiNest with the multimode exploration setting turned on.
+Given these limitations, we instead attempted runs with the more restricted (and hence more computationally tractable) CS- and $P$-informed priors using MultiNest with the multimode exploration setting turned on.
 However, for the CS-informed prior using 4k live points and a sampling efficiency of 0.3, this resulted in seven distinct modes.
 One mode resembled the EOS-agnostic solution, while the six remaining modes were similar to each other in terms of most, but not all, parameters of the CS-informed run with the multimode exploration turned off.
-By contrast, the PP-informed prior proved more tractable.
+By contrast, the $P$-informed prior proved more tractable.
 With 8k live points and a sampling efficiency of 0.3, the multimode run identified two modes: one mode similar to the EOS-agnostic result, and one similar to the EOS-informed prior run, consistent within 95% credible level for all parameters, and within 68% credible level for most parameters.
@@
 (2024).
-Among the EOS-informed prior runs, the CS-informed run is preferred over the PP-informed run according to the Bayes factor, suggesting that the data likely favor smaller radii even for the new geometric mode, consistent with the EOS-agnostic result.
+Among the EOS-informed prior runs, the CS-informed run is preferred over the $P$-informed run according to the Bayes factor, suggesting that the data likely favor smaller radii even for the new geometric mode, consistent with the EOS-agnostic result.
 Nonetheless, serious questions remain about whether the new, more extreme geometric configuration is physical, in particular with regard to the very small (secondary superseding) hot spot.
@@
 To model the EOS MMRR prior for the PPM step, we used a normalizing flow.
-We observed that the normalizing flow struggles to capture the MMRR distribution near $\approx 2.6$ $M_{\odot}$ and 13 km under the PP model, as shown in Fig.
+We observed that the normalizing flow struggles to capture the MMRR distribution near $\approx 2.6$ $M_{\odot}$ and 13 km under the $P$ model, as shown in Fig.
 2.
@@
 The use of EOS priors in the PPM stage results in more stringent constraints on the inferred MRM–R posteriors of PSR J0740+6620 and PSR J0437-4715.
-Including these posteriors in the EOS inference further tightens the PP (CS) EOS posteriors.
-In the MRM–R plane, it produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure-energy density relation for PP (CS), reflecting the expected behavior of these extensions, which relatively favor somewhat stiffer (softer) EOSs during the PPM inference.
+Including these posteriors in the EOS inference further tightens the $P$ (CS) EOS posteriors.
+In the MRM–R plane, it produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure-energy density relation for $P$ (CS), reflecting the expected behavior of these extensions, which relatively favor somewhat stiffer (softer) EOSs during the PPM inference.
 In particular, PSR J0437-4715 has the strongest impact on this effect.
 Studying the impact of EOS-informed PPM results on intermediate-density EOS inferences shows that, for the CS model, the pressures across all three densities consistently shift toward lower values.
-In contrast, for the PP model the pressures shift to higher values across all three densities.
-Moreover, at $3n_{0}$ and $4n_{0}$, the PP-informed pressure posteriors are more peaked than in the EOS-agnostic case.
-Overall, pressures increase for the PP model and decrease for the CS model at all densities, consistent with stiffening and softening of the EOS, respectively, when EOS-informed PPM data are used.
+In contrast, for the $P$ model the pressures shift to higher values across all three densities.
+Moreover, at $3n_{0}$ and $4n_{0}$, the $P$-informed pressure posteriors are more peaked than in the EOS-agnostic case.
+Overall, pressures increase for the $P$ model and decrease for the CS model at all densities, consistent with stiffening and softening of the EOS, respectively, when EOS-informed PPM data are used.
 It is interesting to note that in the EOS analysis with all NICER sources and including gravitational-wave constraints, the prior differences are significantly smaller Rutherford et al.
@@
 865768 AEONS (PI: Watts).
-The work of M.M., I.S., A.S., and K.H.
+The work of $M.$, I.S., A.S., and K.H.
 was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No.